### v1

# Tight Bound for Quantum Unitary Time-Reversal

Abstract
Time-reversal of unitary evolution is fundamental in quantum information processing. Many scenarios, particularly those in quantum learning and metrology, assume free access to the time-reverse of an unknown unitary. In this paper, we settle the query complexity of the unitary time-reversal task: approximately implementing $U^{-1}$ given only black-box access to an unknown $d$-dimensional unitary $U$. We provide a tight query lower bound $\Omega((1-\epsilon)d^{2})$ for the unitary time-reversal to within diamond norm error $\epsilon$. Notably, our lower bound applies to general coherent protocols with unbounded ancillas, and holds even when $\epsilon$ is an average-case distance error. Moreover, our result implies a query lower bound $\Omega(d^{2})$ for approximately implementing control-$U$ up to an irrelevant phase, which is also tight with respect to the dimension.
1 Introduction
The time evolution of a closed quantum system is governed by a reversible unitary $U$. Time-reversal of this evolution can, in principle, be simulated by implementing the inverse unitary $U^{-1}$. However, this requires complete knowledge of the quantum system, which is typically inaccessible when the dynamics is dictated by nature. A straightforward solution is to perform unitary tomography to obtain an approximate description of the unknown $d$-dimensional unitary $U$, which enables approximate time-reversal with error $\epsilon$, using $O(d^{2}/\epsilon)$ queries [HKOT23] to $U$.
In this paper, we consider the unitary time-reversal task: approximately implementing $U^{-1}$ given black-box access to an unknown $d$-dimensional unitary $U$. Beyond its root as a natural concept in physics, this task is fundamental to quantum information theory and has a deep connection to the power of out-of-time-order correlators (OTOCs) [CSM23, SNC+23, SS14, YGS+16, SBSSH16, VES+19, XS24] and advanced quantum learning algorithms with time-reversed unitary access [vACGN23, WZC+22, GP22, MW16, SHH25, Zha25].
Quantum protocols for unitary time-reversal have been extensively studied in the literature [SCHL16, SBZ19, QDS+19a, QDS+19b, EHM+23, QE22, Nav18, GO24, TDN20, TDN23, SST+23, YH21, MZC+25, ZMCW24, MLW25, YSM23, CML+24]. Surprisingly, Yoshida, Soeda, and Murao [YSM23] demonstrated that, when the dimension $d=2$, the unitary time-reversal can be done deterministically and exactly [^1] Exactness and determinism mean that the protocol completes the task without error and with success probability $1$. using only four calls to the unknown unitary. Thereafter, Chen, Mo, Liu, Zhang, and Wang [CML+24] generalized this result by providing a deterministic and exact time-reversal protocol for unitaries of any dimension $d$, using $O(d^{2})$ queries.
Alternatively, if one has additional access to the control-$U$, the unitary time-reversal task can be solved with only $O(\mathrm{poly}(1/\epsilon))$ queries with diamond norm error $\epsilon$, through a folklore use of the phase estimation [CEMM98]. It is worth noting that, this approach uses only a dimension-independent number of queries, though it is intrinsically non-exact. One might naturally wonder whether there exists a (possibly non-exact) protocol that has access only to $U$ and achieves better efficiency than $O(d^{2})$. Hence, the following question remains open:
Is it possible to achieve (approximate) unitary time-reversal using $o(d^{2})$ queries?
Figure 1: Is it possible to achieve (approximate) unitary time-reversal using $o(d^{2})$ queries?
1.1 Our results
In this paper, we provide a negative answer to the above question by showing a tight query lower bound $\Omega((1-\epsilon)d^{2})$. Therefore, our lower bound, combined with the upper bound $O(d^{2})$ by the exact unitary time-reversal protocol given in [CML+24], fully settles the query complexity of the unitary time-reversal task. Formally, our main result is as follows:
Theorem 1.1 (Theorem 3.1 and Corollary 3.2 restated).
Any protocol that solves the unitary time-reversal task to within diamond norm or average-case distance [^2] The average-case distance is given in Definition 2.2. error $\epsilon$, must use at least $\Omega((1-\epsilon)d^{2})$ queries to the unknown $d$-dimensional unitary.
One interesting aspect of our result is the error scaling $(1-\epsilon)$, which fundamentally differs from the typical scaling $\operatorname{poly}(1/\epsilon)$ of learning or metrology tasks. This error scaling is natural as the unitary time-reversal can be done exactly (i.e., $\epsilon=0$). Thus, Theorem 1.1 provides a robust and optimal hardness guarantee: allowing (even to constant-level) error cannot lead to better efficiency. In comparison, most prior works focus on the lower bounds only for the exact case (i.e., $\epsilon=0$), which has recently been settled by Odake, Yoshida, and Murao [OYM24] through a differentiation-based SDP framework (see further comparison in Sections 1.2.1 and 1). We note that their methods cannot provide hardness guarantee in the non-exact regime: it cannot rule out the possibility that a non-exact protocol may achieve better efficiency. It is also worth mentioning that our lower bound holds not only for diamond norm error $\epsilon$ (i.e., worst-case distinguishability), but also for an average-case distance error $\epsilon$ (i.e., average-case distinguishability over Haar random input states), where the latter generally represents a much stronger form of hardness.
Overview of techniques.
Here, we briefly outline the main idea for the proof of Theorem 1.1. Suppose $R$ is a protocol for the unitary time-reversal task that uses $n$ queries to the unknown unitary $U$. Then, $R$ can be described by an $(n+1)$-comb [CDP08, CDP09] on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n+1})$ where each $\mathcal{H}_{i}$ is a copy of $\mathbb{C}^{d}$. We make an additional query to $U$ at the end of $R$, so that the overall channel approximates the identity channel $\mathcal{I}_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}$ from $\mathcal{H}_{0}$ to $\mathcal{H}_{2n+2}$ (see Figure 4). Here, the $n+1$ queries to $U$ in total can be also described by an $(n+1)$-comb:
$C_{U}\coloneqq|U\rangle\!\rangle\!\langle\!\langle U|^{\otimes n+1}.$
Consequently, the overall channel is $R\star C_{U}\approx|I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}|,$ where $\star$ is the link product [CDP09] for quantum combs and $|I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}|$ is the Choi representation of the identity channel $\mathcal{I}_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}$. Taking Haar expectation over $U\in\mathbb{U}_{d}$ and using the bilinearity of $\star$, we obtain:
$R\star\mathop{\bf E\/}_{U}[C_{U}]\approx|I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}|,$ (1)
where $\mathop{\bf E\/}_{U}[C_{U}]$ is the $(n+1)$-th moment of the Haar random unitary under Choi representation (also called performance operator [QE22] for the time-reversal task). Then, we define the stair operators $\{A_{k}\}_{k=1}^{n}$ (see Definition 3.7) which are used to analyze the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$. The stair operators are defined in the Young-Yamanouchi basis [CSST10, Har05] under the Schur-Weyl duality of an asymmetric bipartite quantum system. We prove two lemmas (see Lemma 3.8 and Lemma 3.9) for the stair operators. The first lemma shows that they provide a good upper bound (w.r.t. the Löwner order) for the Haar moment:
$\mathop{\bf E\/}_{U}[C_{U}]\sqsubseteq I_{\mathcal{H}_{2n+1}}\otimes A_{n},$ (2)
where $I_{\mathcal{H}_{2n+1}}$ is the identity operator on $\mathcal{H}_{2n+1}$. The second lemma indicates that the stair operators are compatible with quantum combs through the link product $\star$, i.e., they can be sequentially contracted to identity up to certain coefficient when linked with an arbitrary $(n+1)$-comb $R$:
$R\star(I_{\mathcal{H}_{2n+1}}\otimes A_{n})=\frac{n+1}{d}I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{0}}.$ (3)
Using these two lemmas, we can show
$R\star\mathop{\bf E\/}_{U}[C_{U}]\sqsubseteq R\star(I_{\mathcal{H}_{2n+1}}\otimes A_{n})=\frac{n+1}{d}I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{0}},$ (4)
where we use the property that $\star$ preserves the Löwner order. Intuitively, this means if $n$ is small and $U$ is Haar random, the overall channel $R\star\mathop{\bf E\/}_{U}[C_{U}]$ is highly depolarizing, for any protocol $R$. Therefore, combining Equation 4 with Equation 1, we can see that $n$ must be $\Omega(d^{2})$ since the maximum eigenvalue of $|I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{H}_{0}\rightarrow\mathcal{H}_{2n+2}}|$ is $d$. As our proof is based on the Choi representation, it is natural to expect that our lower bound applies even for the average-case distance error, according to Equation 7.
To prove these two lemmas, we perform a representation-theoretic analysis of the stair operators. Specifically, we exploit the raising and lowering techniques of Young diagrams based on the branching rule of symmetric group representations to show Equation 2 and Equation 3, respectively. Since the subsystem to be traced out is not the “last” subsystem (w.r.t. the ordering fixed by the action of symmetric groups) in the stair operators, the lowering techniques are not directly applicable here. To address this, we use the Young’s orthogonal form [CSST10] for adjacent transpositions to swap the last two subsystems and then analyze the resulting expressions. By this, we can reduce the original problems to pure combinatorics on Young diagrams (see Section 4.2), which is then solved with the assistance of Kerov’s interlacing sequences [Ker93, Ker00].
Lower bound for generalized time-reversal.
Moreover, we can provide a lower bound for the generalized unitary time-reversal task: approximately implement the time-reverse $U^{-t}=e^{-iHt}$ for a given reversing time $t>0$. Here, $H$ is allowed to be any Hamiltonian satisfying $e^{iH}=U$ and $\left\lVert H\right\rVert\leq\pi$. Using Theorem 1.1, we can provide a tight query lower bound $\Omega(d^{2})$ for this task (matching the upper bound $O(d^{2})$ by unitary tomography [HKOT23]). Specifically, we have:
Corollary 1.2 (Corollary 3.3 restated).
Any protocol that that solves the generalized time-reversal task, even for a constant reversing time $t>0$ and to within constant diamond norm error, must use at least $\Omega(d^{2})$ queries to the unknown $d$-dimensional unitary.
Lower bound for unitary controlization.
Our main result can be further used to derive a tight query lower bound for quantum unitary controlization. Specifically, we consider the task called phase-oblivious unitary controlization, i.e., adding quantum control to an unknown unitary $U$ with an unspecified relative phase (see Definition 5.1 for formal definition and discussion there about related work [AFCB14, GST24, TMM+25] on unitary controlization). Since the global phase of $U$ has no physical effect, the task only requires to implement control-$U$ up to an arbitrary relative phase. More specifically, it aims to implement $m$ unitaries:
$V_{1},\ldots,V_{m}\approx\lvert 0\rangle\!\langle 0\rvert\otimes I+e^{i\theta}\lvert 1\rangle\!\langle 1\rvert\otimes U,$
for some $\theta$ not specified. Note that $V_{1},\ldots,V_{m}$ approximate the control-$U$ up to the same relative phase $\theta$. From an operational point of view, this is the minimum requirement for the implemented unitaries $\{V_{i}\}_{i=1}^{m}$ to be then used in quantum algorithms like phase estimation [CEMM98], quantum signal processing [LC17], or quantum singular value transformation [GSLW19]. For this task, we prove a tight query lower bound (matching the upper bound $O(d^{2})$ by unitary tomography [HKOT23]) as follows:
Theorem 1.3 (Theorem 5.2 restated).
Any protocol that solves the phase-oblivious unitary controlization task, even for constant $m$ and to within constant diamond norm error, must use at least $\Omega(d^{2})$ queries to the unknown $d$-dimensional unitary.
The hardness of unitary controlization in Theorem 1.3 is essentially based on the hardness of unitary time-reversal in Theorem 1.1. In particular, by a folklore use of the coherent phase estimation algorithm [CEMM98, NC10] together with a constant number of queries to control-$U$ (with unspecified relative phase), one can implement $U^{-1}$ to within constant error (see Lemma 5.3 for formal statement). This observation, combined with the lower bound for unitary time-reversal, naturally yields a lower bound for unitary controlization.
1.2 Prior works
{tblr}
[…]
($\epsilon=0$) Probabilistic $\Omega(d)$ [QDS+19a, QDS+19b]
Deterministic $\Omega(d^{2})$ [OYM24]
[…]
$\Omega(d)$[YKS+24]
$\Omega(d^{2})$ Our Theorem 1.1
Table 1: Comparison of lower bounds for unitary time-reversal. Here, for non-exact protocols, the lower bounds hold for constant error $\epsilon\geq 0$. Probabilistic exact protocols refers to protocols that are exact conditioned on a flag output bit being “success”.
1.2.1 Lower bounds from semidefinite programming
Quintino, Dong, Shimbo, Soeda, and Murao [QDS+19a, QDS+19b] developed a systematic semidefinite programming (SDP) approach to find probabilistic exact protocols (i.e., exact conditioned on a flag output bit being “success”) for unitary transformation tasks such as transposition, complex conjugation and time-reversal. Query lower bounds are established for such protocols; in particular, $\Omega(d)$ queries are required for probabilistic exact unitary time-reversal. This approach is extended by Yoshida, Koizumi, Studziński, Quintino, and Murao [YKS+24] to derive the same $\Omega(d)$ query lower bound but applicable to approximate unitary time-reversal. Odake, Yoshida, and Murao [OYM24] further developed a general framework for deriving lower bounds on deterministic exact unitary transformation tasks. Specifically, they consider transformations described by differentiable functions $f:\mathbb{U}_{d}\rightarrow\mathbb{U}_{d}$ and analyze the induced Lie algebra mappings obtained by differentiation at various points. Within the Lie algebra framework, they formulate an SDP such that the number of queries required to implement $f(\cdot)$ is lower bounded by the SDP solution. Consequently, they show an $\Omega(d^{2})$ query lower bound for deterministic exact unitary time-reversal, matching the upper bound $O(d^{2})$ by the protocol in [CML+24].
Comparison to [OYM24].
The differentiation-based method in [OYM24] appears challenging to extend directly to the analysis of non-exact protocols, so that their lower bound $\Omega(d^{2})$ only applies to exact unitary time-reversal protocols. This leaves open the possibility that non-exact (approximate) protocols may achieve better efficiency. Indeed, this possibility comes true when access to control-$U$ is available: a non-exact protocol based on phase estimation can approximately implement $U^{-1}$ to within error $\epsilon$ using only $O(\operatorname{poly}(1/\epsilon))$ queries (see Lemma 5.3). One might wonder if this better efficiency is also achievable with access only to $U$. Our lower bound provides a negative answer to this question. Furthermore, as our lower bound applies to non-exact protocols, it can go through the reduction by phase estimation (inherently non-exact) to yield a tight lower bound for unitary contorlization (see Theorem 5.2).
1.2.2 Lower bounds from topological obstructions
Gavorová, Seidel, and Touati [GST24] established lower bounds for unitary transformation tasks from a topological perspective. By proving any $m$-homogeneous function from $\mathbb{U}_{d}$ to the $1$-sphere $\mathcal{S}^{1}$ must satisfy $d\mid m$ (i.e., $d$ divides $m$), they show that any clean protocol (i.e., the ancilla output must be returned in $\lvert 0\rangle$ assuming the protocol is only unitary with postselection) for unitary time-reversal task requires $\Omega(d)$ queries. Note that the cleanness requirement is a strong constraint, since uncomputation of ancillas is generally hard when the protocol has access only to the time-forward $U$. Moreover, they also show that the implementation of $\lvert 0\rangle\!\langle 0\rvert\otimes I+e^{i\theta(U)}\lvert 1\rangle\!\langle 1\rvert\otimes U$ for some pre-fixed phase function $\theta:\mathbb{U}_{d}\rightarrow\mathbb{R}$ cannot be done using finite number of queries to $U$. Note that this task has stronger requirement than our phase-oblivious unitary controlization task (see Definition 5.1) as we do not specify a pre-fixed phase function.
1.2.3 Lower bounds from quantum learning
Cotler, Schuster, and Mohseni [CSM23] designed a quantum learning task such that when ancillas are not allowed: $\Omega(d^{1/4})$ queries to $U$ are required to solve it; and only $O(1)$ queries are sufficient given additional access to the time-reverse $U^{-1}$. This separation implies an $\Omega(d^{1/4})$ query lower bound for the unitary time-reversal task, but it applies only to protocols without ancillas [SHH25]. Schuster, Haferkamp, and Huang [SHH25] later established a stronger separation without this limitation. Their new learning task leverages the low depth construction of pseudorandom unitaries [SHH25, MH25] and the connectivity features of quantum dynamics [SNC+23]. Specifically, this task cannot be solved within any polynomial number of queries to $U$; but it can be solved with $O(1)$ queries given additional access to the time-reverse $U^{-1}$. This implies a lower bound of $\omega(\operatorname{poly}\log d)$ for unitary time-reversal.
1.3 Discussion
In this paper, we settle the complexity of the unitary time-reversal, a fundamental task in quantum information processing. We prove a tight query lower bound $\Omega((1-\epsilon)d^{2})$ for this task. Moreover, our result implies a lower bound $\Omega(d^{2})$ for the unitary contorlization task, which is also tight with respect to the dimension. Here, we further discuss our results from three different aspects.
Hardness for low-depth unitaries.
[…]
General framework for unitary transformations.
An interesting question is how to develop a general framework for the hardness of non-exact unitary transformations, like the SDP-based framework established by [OYM24] for the exact cases. In particular, what are the lower bounds for non-exact unitary transposition and complex conjugation? We conjecture that they both are $\Omega((1-\epsilon)d)$, given the existing lower bound $\Omega(d)$ for exact unitary transposition and complex conjugation by [OYM24, QDS+19a] and the upper bound $O(d)$ for unitary complex conjugation by [MSM19].
Bound for high-precision unitary controlization
In this paper, we provide a query lower bound $\Omega(d^{2})$ for the unitary controlization which is tight with respect to the dimension. However, when considering non-constant error, the current best upper bound is $O(d^{2}/\epsilon)$ (for diamond norm error $\epsilon$), which is obtained from the optimal unitary tomography [HKOT23]. Therefore, a question remains open: Is tomography optimal for this task when concerning non-constant error? Furthermore, can we establish a tight lower bound even for an average-case distance error $\epsilon$?
2 Preliminaries
2.1 Notation
We use $\mathcal{L}(\mathcal{H})$ to denote the set of linear operators on the Hilbert space $\mathcal{H}$. Similarly, we use $\mathcal{L}(\mathcal{H}_{0},\mathcal{H}_{1})$ to denote the set of linear operators from $\mathcal{H}_{0}$ to $\mathcal{H}_{1}$. Given two orthonormal bases for $\mathcal{H}_{0}$ and $\mathcal{H}_{1}$ respectively, we can represent each linear operator in $\mathcal{L}(\mathcal{H}_{0},\mathcal{H}_{1})$ by a $\dim(\mathcal{H}_{1})\times\dim(\mathcal{H}_{0})$ matrix and for such a matrix $X$, we use $|X\rangle\!\rangle\in\mathcal{H}_{1}\otimes\mathcal{H}_{0}$ to denote the vector obtained by flattening the matrix $X$ in row-major order. It is easy to see the following facts:
$|\lvert\psi\rangle\!\langle\phi\rvert\rangle\!\rangle=\lvert\psi\rangle\lvert\phi^{*}\rangle,\quad\quad\quad|XYZ\rangle\!\rangle=X\otimes Z^{\textup{T}}|Y\rangle\!\rangle,$
where $\lvert\phi^{*}\rangle$ is the entry-wise complex conjugate of $\lvert\phi\rangle$ w.r.t. to a given orthonormal basis, and $Z^{\textup{T}}$ is the transpose of the matrix $Z$. The inner product can be denoted by $\langle\!\langle{X}|{Y}\rangle\!\rangle=\operatorname{tr}(X^{\dagger}Y)$.
The Choi-Jamiołkowski operator of a quantum channel $\mathcal{E}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{1})$ is defined by:
$\mathfrak{C}(\mathcal{E})=(\mathcal{E}\otimes\mathcal{I}_{\mathcal{H}_{0}})(|I\rangle\!\rangle\!\langle\!\langle I|),$
where $\mathcal{I}_{\mathcal{H}_{0}}:\mathcal{L}(\mathcal{H}_{0})\rightarrow\mathcal{L}(\mathcal{H}_{0})$ is the identity channel and $|I\rangle\!\rangle=\sum_{i}\lvert i\rangle\lvert i\rangle$ is the (unnormalized) maximally entangled state. Then, the application of the channel $\mathcal{E}$ can be described by its Choi-Jamiołkowski operator:
$\mathcal{E}(X)=\operatorname{tr}_{\mathcal{H}_{0}}\!\left(\mathfrak{C}(\mathcal{E})^{\textup{T}_{\mathcal{H}_{0}}}\cdot(I_{\mathcal{H}_{1}}\otimes X)\right),$
where $\operatorname{tr}_{\mathcal{H}_{0}}$ and $\textup{T}_{\mathcal{H}_{0}}$ denote the partial trace and partial transpose on the subsystem $\mathcal{H}_{0}$, respectively.
For a sequence of Hilbert spaces $\mathcal{H}_{0},\ldots,\mathcal{H}_{n}$ indexed by consecutive integers, we use $\mathcal{H}_{i:j}$ to denote the Hilbert space $\mathcal{H}_{i}\otimes\cdots\otimes\mathcal{H}_{j}$. For two linear operators $X,Y$, we use $X\sqsubseteq Y$ to denote that $Y-X$ is positive semidefinite.
2.2 Distance between quantum channels
Worst-case distance.
First, we introduce the diamond norm, which servers as the worst-case distance between quantum channels.
Definition 2.1 (Diamond norm [AKN98]).
The diamond norm of two quantum channels $\mathcal{E}_{1},\mathcal{E}_{2}$ is defined as
$\|\mathcal{E}_{1}-\mathcal{E}_{2}\|_{\diamond}\coloneqq\sup_{\rho}\|(\mathcal{E}_{1}\otimes\mathcal{I})(\rho)-(\mathcal{E}_{2}\otimes\mathcal{I})(\rho)\|_{1},$
where $\|\cdot\|_{1}$ denotes the Schatten $1$-norm (trace norm).
There is a simple relation between the diamond norm of quantum channels and trace norm of the corresponding Choi-Jamiołkowski operators. Specifically, let $\mathcal{E}_{1},\mathcal{E}_{2}:\mathcal{L}(\mathbb{C}^{d})\rightarrow\mathcal{L}(\mathbb{C}^{d})$ be two $d$-dimensional quantum channels. Then, by the definition of diamond norm, we have
$\|\mathcal{E}_{1}-\mathcal{E}_{2}\|_{\diamond}$ $\geq\left\|(\mathcal{E}_{1}\otimes\mathcal{I})\!\left(\frac{|I\rangle\!\rangle\!\langle\!\langle I|}{d}\right)-(\mathcal{E}_{2}\otimes\mathcal{I})\!\left(\frac{|I\rangle\!\rangle\!\langle\!\langle I|}{d}\right)\right\|_{1}$
$=\frac{1}{d}\|\mathfrak{C}(\mathcal{E}_{1})-\mathfrak{C}(\mathcal{E}_{2})\|_{1},$ (5)
where $|I\rangle\!\rangle\!\langle\!\langle I|/d$ is a maximally entangled state.
Average-case distance.
Then, we introduce the average-case distance between quantum channels, which is widely used in quantum learning [HLB+24, VH25], benchmarking [MGE12, MGE11] and metrology [CWLY23, YRC20]. For this, we first need the notion of fidelity. Recall that the fidelity of two quantum states $\rho,\sigma$ is defined as
$F(\rho,\sigma)\coloneqq\operatorname{tr}\!\left(\sqrt{\sqrt{\sigma}\rho\sqrt{\sigma}}\right)^{2}.$
Note that when $\sigma=\lvert\psi\rangle\!\langle\psi\rvert$ is a pure state, we have $F(\rho,\sigma)=\operatorname{tr}(\rho\sigma)=\langle\psi\rvert\rho\lvert\psi\rangle$. Then, the definition of average-case distance is as follows:
Definition 2.2 (Average-case distance [Nie02, HLB+24, VH25]).
The average-case distance between two quantum channels $\mathcal{E}_{1},\mathcal{E}_{2}$ is defined as
$\mathcal{D}_{\textup{avg}}(\mathcal{E}_{1},\mathcal{E}_{2})\coloneqq\mathop{\bf E\/}_{\lvert\psi\rangle}[1-F(\mathcal{E}_{1}(\lvert\psi\rangle\!\langle\psi\rvert),\mathcal{E}_{2}(\lvert\psi\rangle\!\langle\psi\rvert))],$
where $\mathop{\bf E\/}_{\lvert\psi\rangle}$ denotes the expectation over the Haar random state $\lvert\psi\rangle$.
First, it is easy to see that the average-case distance is unitarily invariant, i.e.,
$\mathcal{D}_{\textup{avg}}(\mathcal{U}\circ\mathcal{E}_{1}\circ\mathcal{V},\,\mathcal{U}\circ\mathcal{E}_{2}\circ\mathcal{V})=\mathcal{D}_{\textup{avg}}(\mathcal{E}_{1},\mathcal{E}_{2}),$ (6)
for any unitary channels $\mathcal{U},\mathcal{V}$. Another property is that, when one of $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$ is a unitary channel, this distance simplifies to the average gate fidelity and is closely related to the entanglement gate fidelity [Nie02]. Specifically, we have
$\mathcal{D}_{\textup{avg}}(\mathcal{E},\mathcal{U})=\frac{d}{d+1}\left(1-F\!\left(\frac{\mathfrak{C}(\mathcal{E})}{d},\frac{|U\rangle\!\rangle\!\langle\!\langle U|}{d}\right)\right),$ (7)
where $d$ is the dimension of $\mathcal{U}$ and $\mathcal{E}$, $F(\mathfrak{C}(\mathcal{E})/d,|U\rangle\!\rangle\!\langle\!\langle U|/d)=\operatorname{tr}(|U\rangle\!\rangle\!\langle\!\langle U|\cdot\mathfrak{C}(\mathcal{E}))/d^{2}$ is the entanglement gate fidelity between $\mathcal{U}$ and $\mathcal{E}$. Then, it is easy to see that
$\mathcal{D}_{\textup{avg}}(\mathcal{E},\mathcal{U})$ $\leq\frac{d}{d+1}\left\|\frac{\mathfrak{C}(\mathcal{E})}{d}-\frac{|U\rangle\!\rangle\!\langle\!\langle U|}{d}\right\|_{1}\leq\frac{d}{d+1}\|\mathcal{E}-\mathcal{U}\|_{\diamond},$ (8)
where the first inequality is due to the well-known property [FVDG02]: $1-\sqrt{F(\rho,\sigma)}\leq\frac{1}{2}\|\rho-\sigma\|_{1}$ and $F(\rho,\sigma)\leq 1$, the second inequality is by Equation 5.
We remark that the average-case distance can be much smaller than the diamond norm. For example, let $U=(\lvert 0\rangle\!\langle 1\rvert+\lvert 1\rangle\!\langle 0\rvert)\oplus I_{d-2}$ be a $d$-dimensional unitary, where $I_{d-2}$ is the identity operator of dimension $d-2$. In such case, we have $\mathcal{D}_{\textup{avg}}(\mathcal{U},\mathcal{I})=O(1/d)$ but $\|\mathcal{U}-\mathcal{I}\|_{\diamond}=\Omega(1)$. This implies that hardness results with average-case distance are generally much stronger than those with diamond norm.
2.3 Quantum combs
In this section, we introduce the quantum comb [CDP08, CDP09], which is a powerful tool to describe (higher) transformations of quantum processes. Specifically, the Choi-Jamiołkowski representation of quantum channels (i.e., transformations of quantum states) can be generalized to a higher-level concept (i.e., transformations of quantum processes), which is called quantum comb.
Definition 2.3 (Quantum comb [CDP09]).
For an integer $n\geq 1$, a quantum $n$-comb defined on a sequence of $2n$ Hilbert spaces $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n-1})$ is a positive semidefinite operator $X$ on $\bigotimes_{j=0}^{2n-1}\mathcal{H}_{j}$ such that there exists a sequence of operators $X^{(n)},X^{(n-1)},\ldots,X^{(1)},X^{(0)}$ such that
$\begin{split}\operatorname{tr}_{\mathcal{H}_{2j-1}}\!\left(X^{(j)}\right)&=I_{\mathcal{H}_{2j-2}}\otimes X^{(j-1)},\quad 1\leq j\leq n,\end{split}$ (9)
where $X^{(n)}=X$ and $X^{(0)}=1$.
Note that in Definition 2.3, the operators $X^{(j)}$ are again quantum $j$-combs (for convenience, we define any quantum 00-comb to be the scalar $1$). It is also worth noting that a quantum $1$-comb is simply the Choi-Jamiołkowski operator of a quantum channel.
Figure 2: The combination of a $4$-comb $X$ with a $3$-comb $Y$, resulting in a $1$-comb $X\star Y$ on $(\mathcal{H}_{0},\mathcal{H}_{7})$.
The combination of any two quantum combs is defined by the link product $\star$”:
Definition 2.4 (Link product $\star$” [CDP09]).
Suppose $X$ is a linear operator on $\mathcal{H}_{\bm{i}}=\mathcal{H}_{i_{1}}\otimes\mathcal{H}_{i_{2}}\otimes\cdots\otimes\mathcal{H}_{i_{n}}$ and $Y$ is a linear operator on $\mathcal{H}_{\bm{j}}=\mathcal{H}_{j_{1}}\otimes\mathcal{H}_{j_{2}}\otimes\cdots\otimes\mathcal{H}_{j_{m}}$, where $\bm{i}=(i_{1},\ldots,i_{n})$ is a sequence of pairwise distinct indices, and similar for $\bm{j}=(j_{1},\ldots,j_{m})$. Let $\bm{a}=\bm{i}\cap\bm{j}$ be the set of indices in both $\bm{i}$ and $\bm{j}$ and $\bm{b}=\bm{i}\cup\bm{j}$ be the set of indices in either $\bm{i}$ or $\bm{j}$. Then, the combination of $X$ and $Y$ is defined by
$X\star Y=\operatorname{tr}_{\mathcal{H}_{\bm{a}}}\!\left(X^{\textup{T}_{\mathcal{H}_{\bm{a}}}}\cdot Y\right)=\operatorname{tr}_{\mathcal{H}_{\bm{a}}}\!\left(X\cdot Y^{\textup{T}_{\mathcal{H}_{{\bm{a}}}}}\right),$
where $\mathcal{H}_{\bm{a}}$ means the tensor product of subsystems labeled by the indices in $\bm{a}$, $\textup{T}_{\mathcal{H}_{\bm{a}}}$ means the partial transpose on $\mathcal{H}_{\bm{a}}$, both $X$ and $Y$ are treated as linear operators on $\mathcal{H}_{\bm{b}}$, extended by tensoring with the identity operator as needed.
The link product has a very good property [CDP09, Theorem 2]:
$X,Y\sqsupseteq 0\Longrightarrow X\star Y\sqsupseteq 0.$ (10)
Moreover, it characterizes the channel concatenation under the Choi representation: given two quantum channels $\mathcal{E}_{1}:\mathcal{L}(\mathcal{H}_{1})\rightarrow\mathcal{L}(\mathcal{H}_{2})$ and $\mathcal{E}_{2}:\mathcal{L}(\mathcal{H}_{2})\rightarrow\mathcal{L}(\mathcal{H}_{3})$, we have $\mathfrak{C}(\mathcal{E}_{2}\circ\mathcal{E}_{1})=\mathfrak{C}(\mathcal{E}_{2})\star\mathfrak{C}(\mathcal{E}_{1}).$ The link product is also used to describe the combination of two higher-order combs such as that shown in Figure 2. More generally, suppose $X$ is an $n$-comb on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n-1})$ and $Y$ is an $(n-1)$-comb on $(\mathcal{H}_{1},\mathcal{H}_{2}\,\ldots,\mathcal{H}_{2n-2})$, then
$X\star Y=\operatorname{tr}_{\mathcal{H}_{1:2n-2}}\!\left(X^{\textup{T}_{\mathcal{H}_{1:2n-2}}}\cdot(I_{\mathcal{H}_{2n-1}}\otimes Y\otimes I_{\mathcal{H}_{0}})\right)=\operatorname{tr}_{\mathcal{H}_{1:2n-2}}\!\left(X\cdot(I_{\mathcal{H}_{2n-1}}\otimes Y^{\textup{T}}\otimes I_{\mathcal{H}_{0}})\right)$
turns out to be a $1$-comb on $(\mathcal{H}_{0},\mathcal{H}_{2n-1})$. To see this, first note that $X,Y$ are positive semidefinite, which means $X\star Y$ is also positive semidefinite. On the other hand, note that
$\operatorname{tr}_{\mathcal{H}_{2n-1}}(X\star Y)$ $=\operatorname{tr}_{\mathcal{H}_{1:2n-2}}\!\left(\operatorname{tr}_{\mathcal{H}_{2n-1}}(X)^{\textup{T}_{\mathcal{H}_{1:2n-2}}}\cdot(Y\otimes I_{\mathcal{H}_{0}})\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2n-2}}\!\left((I_{\mathcal{H}_{2n-2}}\otimes X^{(n-1)})^{\textup{T}_{\mathcal{H}_{1:2n-2}}}\cdot(Y\otimes I_{\mathcal{H}_{0}})\right)$ (11)
$=\operatorname{tr}_{\mathcal{H}_{1:2n-3}}\!\left((X^{(n-1)})^{\textup{T}_{\mathcal{H}_{1:2n-3}}}\cdot(\operatorname{tr}_{\mathcal{H}_{2n-2}}(Y)\otimes I_{\mathcal{H}_{0}})\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2n-3}}\!\left((X^{(n-1)})^{\textup{T}_{\mathcal{H}_{1:2n-3}}}\cdot(I_{\mathcal{H}_{2n-3}}\otimes Y^{(n-2)}\otimes I_{\mathcal{H}_{0}})\right)$ (12)
$=\operatorname{tr}_{\mathcal{H}_{1:2n-3}}\!\left(((X^{(n-1)})^{\textup{T}_{\mathcal{H}_{1:2n-4}}})^{\textup{T}_{\mathcal{H}_{2n-3}}}\cdot((I_{\mathcal{H}_{2n-3}})^{\textup{T}}\otimes Y^{(n-2)}\otimes I_{\mathcal{H}_{0}})\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2n-3}}\!\left((X^{(n-1)})^{\textup{T}_{\mathcal{H}_{1:2n-4}}}\cdot(I_{\mathcal{H}_{2n-3}}\otimes Y^{(n-2)}\otimes I_{\mathcal{H}_{0}})\right)$ (13)
$=\operatorname{tr}_{\mathcal{H}_{2n-3}}\!\left(X^{(n-1)}\star Y^{(n-2)}\right),$
where in Equation 11, $X^{(n-1)}$ is an $(n-1)$-comb obtained from taking partial trace on $X$ (see Definition 2.3) and similarly in Equation 12, $Y^{(n-2)}$ is an $(n-2)$-comb obtained from $Y$, Equation 13 is due to the property of partial trace $\operatorname{tr}_{\mathcal{H}}(X^{\textup{T}_{\mathcal{H}}})=\operatorname{tr}_{\mathcal{H}}(X)$. Then, we can iterate this process and finally obtain
$\operatorname{tr}_{\mathcal{H}_{2n-1}}(X\star Y)=\operatorname{tr}_{\mathcal{H}_{1}}\!\left(X^{(1)}\right)=I_{\mathcal{H}_{0}},$
which means $X\star Y$ is a $1$-comb on $(\mathcal{H}_{0},\mathcal{H}_{2n-1})$.
2.4 Young diagrams
Young diagrams and standard Young tableaux.
A Young diagram $\lambda$ consisting of $n$ boxes and $\ell$ rows is a partition $(\lambda_{1},\ldots,\lambda_{\ell})$ of $n$ such that $\sum_{i=1}^{\ell}\lambda_{i}=n$ and $\lambda_{1}\geq\cdots\geq\lambda_{\ell}>0$. The $i$-th rows consists of $\lambda_{i}$ boxes. By convention, the Young diagram is drawn with left-justified rows, arranged from top to bottom. The conjugate of a Young diagram is obtained by exchanging rows and columns. For example, the Young diagram $\lambda=(4,3,1)$ and its conjugate $\lambda^{\prime}=(3,2,2,1)$ are drawn as:
{ytableau}
 and
{ytableau}
{ytableau}
 and
{ytableau}
[…]
respectively. We will use $\lambda\vdash n$ to denote that $\lambda$ is a Young diagram with $n$ boxes.
A standard Young tableau of shape $\lambda\vdash n$ is obtained by a filling of the $n$ boxes of $\lambda$ using the integers $1,\ldots,n$, each appearing exactly once, such that the entries increase from left to right in each row and from top to bottom in each column. An example of standard Young tableau of shape $(4,3,1)$ is
{ytableau}
1&346
278
5
{ytableau}
1&346
278
5
[…]
Without causing confusion, we refer to the box in $T$ containing the integer $i$ as the $i$-box. We use $\textup{Sh}(T)$ to denote the shape (Young diagram) of a standard Young tableau $T$. We use $\textup{Tab}(\lambda)$ to denote the set of all standard Young tableaux $T$ such that $\textup{Sh}(T)=\lambda$ and use $\textup{Tab}(n)$ to denote $\bigcup_{\lambda\vdash n}\textup{Tab}(\lambda)$. Given a Young diagram $\lambda\vdash n$, the number of standard Young tableaux of shape $\lambda$ (i.e., $|\textup{Tab}(\lambda)|$) is characterized by the hook length formula [Ful97]:
$|\textup{Tab}(\lambda)|=\frac{n!}{\prod_{\square\in\lambda}h_{\lambda}(\square)},$ (14)
where $h_{\lambda}(\square)$ is the number of boxes that are directly to the right and below the box $\square$, including the box itself (i.e., the “hook length”).
Adjacency relation.
For $\lambda\vdash n+1$, $\mu\vdash n$, we write $\mu\nearrow\lambda$ or $\lambda\searrow\mu$ if $\lambda$ can be obtained from $\mu$ by adding one box. For a standard Young tableau $T$, we use $T^{\downarrow}$ to denote the standard Young tableau obtained from $T$ by removing the box containing the largest integer in $T$. Suppose $\mu\nearrow\lambda$ and $T\in\textup{Tab}(\mu)$. We use $T^{\uparrow\lambda}$ to denote the standard Young tableau obtained from $T$ by adding a box containing number $n+1$, resulting in shape $\lambda$. We define $\textup{Tab}(\lambda,\mu)\subseteq\textup{Tab}(\lambda)$ to be the set of all standard Young tableaux $T$ such that $\textup{Sh}(T)=\lambda$ and $\textup{Sh}(T^{\downarrow})=\mu$. It is easy to see that:
$|\textup{Tab}(\lambda,\mu)|=|\textup{Tab}(\mu)|.$
Let $\mu,\nu\vdash n$ and $\mu\neq\nu$. We call $\mu$ and $\nu$ are adjacent to each other if one of the following equivalent conditions is true:
  • There exists a $\lambda\vdash n+1$ such that $\mu\nearrow\lambda$ and $\nu\nearrow\lambda$.
  • There exists a $\tau\vdash n-1$ such that $\tau\nearrow\mu$ and $\tau\nearrow\nu$.
Such $\mu$ and $\nu$ can be obtained from each other by moving a removable box. Furthermore, $\lambda=\mu\cup\nu$ and $\tau=\mu\cap\nu$ are unique, where $\mu\cup\nu$ denotes the Young diagram consisting of the union of boxes in $\mu$ and in $\nu$, and $\mu\cap\nu$ denotes the Young diagram consisting of the intersection of boxes in $\mu$ and in $\nu$. Moreover, we use $\lambda\setminus\mu$ to denote the set of boxes (not a Young diagram) in $\lambda$ but not in $\mu$.
Kerov’s interlacing sequences.
Kerov’s interlacing sequences [Ker93, Ker00] serve as an important tool in studying the combinatorics of Young diagram and their analytic generalizations. Suppose $\lambda\vdash n$ is a Young diagram. The rotated version of $\lambda$ is defined by rotating the Young diagram $\lambda^{\prime}$ (the conjugate of $\lambda$) $135^{\circ}$ counterclockwise. We define the $x$-axis so that the center of each box lies at an integer $x$-axis coordinate. Consider the shape of the rotated Young diagram extended with the linear curves $y=x$ and $y=-x$ (e.g., the bold line in Figure 3(a)), which is a piecewise linear function. Then, Kerov’s interlacing sequences are defined as the sequence $x$-coordinates of the local minima $\bm{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{L})$ and the sequence of $x$-coordinates of the local maxima $\bm{\beta}=(\beta_{1},\beta_{2},\ldots,\beta_{L-1})$. The two sequences interlace
$\alpha_{1}<\beta_{1}<\alpha_{2}<\cdots<\alpha_{L-1}<\beta_{L-1}<\alpha_{L}.$
Figure 3(a) shows an example of the Young diagram $(6,5,3,3,2,1,1,1)$ and its interlacing sequences are $\bm{\alpha}=(-8,-4,-2,1,4,6)$ and $\bm{\beta}=(-7,-3,-1,3,5)$.
(a) The interlacing sequences of $\lambda$.
(b) Axial coordinates of two boxes in $\lambda$.
Figure 3: Rotated version of the Young diagram $\lambda=(6,5,3,3,2,1,1,1)$.
We define the axial coordinate [^5] The axial coordinate is also called content in the literature [CSST10, OV96] $c(\square)$ for a box $\square$ in a Young diagram as the $x$-axis coordinate of the center of this box when drawing in the rotated version. Figure 3(b) shows an example of the Young diagram $(6,5,3,3,2,1,1,1)$ and the axial coordinates of two boxes. The axial distance from box $\square_{1}$ to box $\square_{2}$ is defined as the difference $c(\square_{1})-c(\square_{2})$.
Adding or removing a box from a Young diagram can be described naturally using Kerov’s interlacing sequences. A position where a box can be added corresponds precisely to a local minima $\alpha_{i}$. Conversely, a removable box corresponds to a local maxima $\beta_{i}$. Specifically, the axial coordinate of the added (or removed) box $\square$ (i.e., $c(\square)$) coincides with the $\alpha_{i}$ (or $\beta_{i}$). Suppose $\mu\vdash n$ has interlacing sequences $\alpha_{1},\ldots,\alpha_{L}$ and $\beta_{1},\ldots,\beta_{L-1}$. Let $\lambda\vdash n+1$ be obtained from $\mu$ by adding a box at $\alpha_{k}$. Then, we have [Ker93, Eq. (3.2.3)]
$\frac{|\textup{Tab}(\lambda)|}{|\textup{Tab}(\mu)|}=(n+1)\prod_{i=1}^{k-1}\frac{\alpha_{k}-\beta_{i}}{\alpha_{k}-\alpha_{i}}\prod_{i=k+1}^{L}\frac{\alpha_{k}-\beta_{i-1}}{\alpha_{k}-\alpha_{i}}.$ (15)
On the other hand, suppose $\lambda\vdash n$ has interlacing sequences $\alpha_{1},\ldots,\alpha_{L}$ and $\beta_{1},\ldots,\beta_{L-1}$. Let $\mu\vdash n-1$ be obtained from $\lambda$ by removing a box at $\beta_{k}$. Then, we have [Ker00, Lemma 3.3]
$\frac{|\textup{Tab}(\mu)|}{|\textup{Tab}(\lambda)|}=\frac{(\alpha_{L}-\beta_{k})(\beta_{k}-\alpha_{1})}{n}\prod_{i=1}^{k-1}\frac{\beta_{k}-\alpha_{i+1}}{\beta_{k}-\beta_{i}}\prod_{i=k+1}^{L-1}\frac{\beta_{k}-\alpha_{i}}{\beta_{k}-\beta_{i}}.$ (16)
2.5 Schur-Weyl duality
Let $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{n}$ be a sequence of Hilbert space such that $\mathcal{H}_{i}\cong\mathbb{C}^{d}$ for $1\leq i\leq n$. Consider the Hilbert space $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$. This space admits representations of the symmetric group $\mathfrak{S}_{n}$ (i.e., the group of all permutations on the set $\{1,2,\ldots,n\}$) and unitary group $\mathbb{U}_{d}$ (i.e., the group of unitaries on $d$-dimensional Hilbert space). The unitary group acts by simultaneous “rotation” as $U^{\otimes n}$ for any $U\in\mathbb{U}_{d}$ and the symmetric group acts by permuting tensor factors:
$P(\pi)\lvert\psi_{1}\rangle\cdots\lvert\psi_{n}\rangle=\lvert\psi_{\pi^{-1}(1)}\rangle\cdots\lvert\psi_{\pi^{-1}(n)}\rangle,$ (17)
where $\pi\in\mathfrak{S}_{n}$.
Remark 2.5.
In this paper, when we say the symmetric group $\mathfrak{S}_{n}$ acts on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$, the order of tensor product matters instead of the indices $i$. For example, if $\mathfrak{S}_{3}$ acts on $\mathcal{H}_{1}\otimes\mathcal{H}_{4}\otimes\mathcal{H}_{2}$, then we have
$P(\pi_{(13)})\lvert\psi_{1}\rangle\lvert\psi_{4}\rangle\lvert\psi_{2}\rangle=\lvert\psi_{2}\rangle\lvert\psi_{4}\rangle\lvert\psi_{1}\rangle,$
for $\lvert\psi_{i}\rangle\in\mathcal{H}_{i}$, where $\pi_{(13)}\in\mathfrak{S}_{3}$ is the permutation that swaps integers $1$ and $3$.
Two actions $U^{\otimes n}$ and $P(\pi)$ commute with each other, and hence $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$ admits a representation of group $\mathbb{U}_{d}\times\mathfrak{S}_{n}$. More specifically, the Schur-Weyl duality (see, e.g., [FH13]) states that
$\bigotimes_{i=1}^{n}\mathcal{H}_{i}\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathfrak{S}_{n}}}{{\cong}}\bigoplus_{\lambda\vdash n}\mathcal{Q}_{\lambda}\otimes\mathcal{P}_{\lambda},$
where $\mathcal{Q}_{\lambda}$ and $\mathcal{P}_{\lambda}$ are irreducible representations of $\mathbb{U}_{d}$ and $\mathfrak{S}_{n}$ labeled by Young diagram $\lambda$, respectively. [^6] Note that $\mathcal{Q}_{\lambda}=0$ if the number of rows of $\lambda$ is greater than $d$.
Now, suppose we have two sequences of Hilbert spaces $(\mathcal{H}^{\textup{A}}_{1},\ldots,\mathcal{H}^{\textup{A}}_{n})$ and $(\mathcal{H}^{\textup{B}}_{1},\ldots,\mathcal{H}^{\textup{B}}_{m})$, where $\mathcal{H}^{\textup{A}}_{i}\cong\mathcal{H}^{\textup{B}}_{j}\cong\mathbb{C}^{d}$. We define the action of group $\mathfrak{S}_{n}\times\mathfrak{S}_{m}$ on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}^{\textup{A}}\otimes\bigotimes_{j=1}^{m}\mathcal{H}_{j}^{\textup{B}}$ as $P(\pi^{\textup{A}})\otimes P(\pi^{\textup{B}})$ for $(\pi^{\textup{A}},\pi^{\textup{B}})\in\mathfrak{S}_{n}\times\mathfrak{S}_{m}$. Similarly, we define the action of $\mathbb{U}_{d}\times\mathbb{U}_{d}$ on $\bigotimes_{i=1}^{n}\mathcal{H}^{\textup{A}}_{i}\otimes\bigotimes_{j=1}^{m}\mathcal{H}^{\textup{B}}_{j}$ as $(U^{\textup{A}})^{\otimes n}\otimes(U^{\textup{B}})^{\otimes m}$ for $(U^{\textup{A}},U^{\textup{B}})\in\mathbb{U}_{d}\times\mathbb{U}_{d}$. Note that the action of $\mathfrak{S}_{n}\times\mathfrak{S}_{m}$ commutes with the action of $\mathbb{U}_{d}\times\mathbb{U}_{d}$. Therefore, we have a generalized Schur-Weyl duality on this bipartite system as
$\bigotimes_{i=1}^{n}\mathcal{H}^{\textup{A}}_{i}\otimes\bigotimes_{j=1}^{m}\mathcal{H}^{\textup{B}}_{j}\,\,\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{n}\times\mathfrak{S}_{m}}}{{\cong}}\,\,\bigoplus_{\begin{subarray}{c}\lambda\vdash n\\ \mu\vdash m\end{subarray}}\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu},$
where $\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu}$ is an irreducible representation of $\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{n}\times\mathfrak{S}_{m}$.
2.6 Young-Yamanouchi basis
Branching rule of $\mathfrak{S}_{n}$.
We now look deeper into the structure of $\mathcal{P}_{\lambda}$. Consider the following chain of groups:
$\mathfrak{S}_{1}\subset\mathfrak{S}_{2}\subset\cdots\subset\mathfrak{S}_{n},$
where the embedding $\mathfrak{S}_{i}\rightarrow\mathfrak{S}_{i+1}$ is by identifying each $\pi\in\mathfrak{S}_{i}$ as the permutation in $\mathfrak{S}_{i+1}$ that fixes the element $i+1$. Note that the irreducible representations of $\mathfrak{S}_{i}$ are in one-to-one correspondence to the Young diagrams consisting of $i$ boxes. Let $\lambda\vdash i$ be a Young diagram of $i$ boxes and $\mathcal{P}_{\lambda}$ be an irreducible representation of $\mathfrak{S}_{i}$. Then, $\mathcal{P}_{\lambda}$ is also a representation of $\mathfrak{S}_{i-1}$, which decomposes as a direct sum of irreducible representations of $\mathfrak{S}_{i-1}$. More specifically, we have the following branching rule [CSST10, Mél17]:
$\textup{Res}^{\mathfrak{S}_{i}}_{\mathfrak{S}_{i-1}}\mathcal{P}_{\lambda}=\bigoplus_{\begin{subarray}{c}\mu:\mu\nearrow\lambda\end{subarray}}\mathcal{P}_{\mu},$ (18)
where $\textup{Res}^{\mathfrak{S}_{i}}_{\mathfrak{S}_{i-1}}$ denotes the restriction of representation from $\mathfrak{S}_{i}$ to $\mathfrak{S}_{i-1}$, $\mu\nearrow\lambda$ means that $\lambda$ can be obtained from $\mu$ by adding one box. Equation 18 means that we can decompose the vector space $\mathcal{P}_{\lambda}$ into direct sum of smaller vector spaces. By iterating this process (i.e., taking further restrictions on $\mathcal{P}_{\mu}$), we finally obtain a direct sum of the trivial one-dimensional representations of $\mathfrak{S}_{1}$. Therefore, this process determines a basis (up to scalar factors) of $\mathcal{P}_{\lambda}$ and each basis vector can be parameterized by its branching path:
$\lambda^{(1)}\rightarrow\lambda^{(2)}\rightarrow\cdots\rightarrow\lambda^{(n)}=\lambda,$ (19)
where $\lambda^{(i)}\nearrow\lambda^{(i+1)}$ and $\lambda^{(1)}$ is the one-box Young diagram. This basis is called Young-Yamanouchi basis or Gelfand-Tsetlin basis for $\mathfrak{S}_{n}$ [OV96]. For convenience, we will simply call it Young basis in this paper.
Furthermore, we can see that the path in Equation 19 one-to-one corresponds to a standard Young tableau $T$ of shape $\lambda=\lambda^{(n)}$, where the integer $i$ is assigned to the box $\lambda^{(i)}\setminus\lambda^{(i-1)}$, in which $\lambda^{(i)}\setminus\lambda^{(i-1)}$ denotes the shape by removing from the shape of $\lambda^{(i)}$ all the boxes belonging to $\lambda^{(i-1)}$. By this mean, we may also denote the standard Young tableau as
$T=\lambda^{(1)}\rightarrow\lambda^{(2)}\rightarrow\cdots\rightarrow\lambda^{(n)}.$
By choosing an appropriate inner product making $\mathcal{P}_{\lambda}$ a unitary representation of $\mathfrak{S}_{n}$ and then normalizing the Young basis w.r.t. the inner product, we thus obtain an orthonormal basis $\{\lvert T\rangle\}_{T\in\textup{Tab}(\lambda)}$ of $\mathcal{P}_{\lambda}$ labeled by standard Young tableaux $T\in\textup{Tab}(\lambda)$. This also implies that $\dim(\mathcal{P}_{\lambda})=|\textup{Tab}(\lambda)|$.
Isotypic component projectors.
Consider the group algebra $\mathbb{C}\mathfrak{S}_{n}$, which is the associative algebra containing formal linear combinations of the elements in $\mathfrak{S}_{n}$ with coefficients in $\mathbb{C}$. Note that any representation of $\mathfrak{S}_{n}$ is naturally a representation of $\mathbb{C}\mathfrak{S}_{n}$ by linearly extending the group action of $\mathfrak{S}_{n}$. For any $1\leq k\leq n$ and $\lambda\vdash k$, we define the following element
$e_{\lambda}\coloneqq\frac{\dim(\mathcal{P}_{\lambda})}{k!}\sum_{\pi\in\mathfrak{S}_{k}}\chi^{*}_{\lambda}(\pi)\pi\in\mathbb{C}\mathfrak{S}_{k}\subseteq\mathbb{C}\mathfrak{S}_{n},$ (20)
where $\chi_{\lambda}(\cdot)$ is the character of $\mathcal{P}_{\lambda}$. By the orthogonality of characters (see e.g., [EGH+11, Sec. 4.5]), $e_{\lambda}$ acts on $\mathcal{P}_{\lambda}$ as the identity and acts on $\mathcal{P}_{\mu}$ as the null map for $\mu\vdash k,\mu\neq\lambda$, thus is the projector onto the isotypic component of $\lambda$ (in fact, for any unitary representation of $\mathfrak{S}_{k}$, $e_{\lambda}$ acts as an orthogonal projector). Therefore, for any standard Young tableau $T=\lambda^{(1)}\rightarrow\cdots\rightarrow\lambda^{(n)}$, we have
$e_{\lambda^{(k)}}\lvert T\rangle=\lvert T\rangle,$ (21)
for any $1\leq k\leq n$. Moreover, $\lvert T\rangle$ is the only vector (up to a scalar) satisfying Equation 21 for any $1\leq k\leq n$. This means, under the algebra isomorphism $\mathbb{C}\mathfrak{S}_{n}\cong\bigoplus_{\mu\vdash n}\mathcal{L}(\mathcal{P}_{\mu})$ (due to the semisimplicity of $\mathbb{C}\mathfrak{S}_{n}$), we can write
$\lvert T\rangle\!\langle T\rvert=\prod_{k=1}^{n}e_{\lambda^{(k)}},$ (22)
where $\lvert T\rangle\!\langle T\rvert$ is considered as a linear operator in $\bigoplus_{\mu\vdash n}\mathcal{L}(\mathcal{P}_{\mu})$.
Young’s orthogonal form.
For $11\leq i\leq n-$ 1, let $s_{i}\in\mathfrak{S}_{n}$ be the $i$-th adjacent transposition (i.e., $s_{i}$ swaps integers $i$ and $i+1$). It is known that those $s_{i}$ generate the group $\mathfrak{S}_{n}$ and are called Coxeter generators. The Young’s orthogonal form [CSST10] gives an explicit description of the action of $s_{i}$ on the Young basis:
$s_{i}\lvert T\rangle=\frac{1}{r(T)}\lvert T\rangle+\sqrt{1-\frac{1}{r(T)^{2}}}\lvert s_{i}T\rangle,$ (23)
where $r(T)$ is the axial distance from the $(i+1)$-box to the $i$-box in $T$, $s_{i}T$ is the Young tableau obtained from $T$ by swapping the $i$-box and the $(i+1)$-box. [^7] Note that $s_{i}T$ may not be a standard Young tableau. Nevertheless, in such case $\sqrt{1-1/r(T)^{2}}=0$.
3 Hardness of unitary time-reversal
In this section, we will prove our lower bounds for quantum unitary time-reversal. First, we provide a tight lower bound for time-reversal with the average-case distance error (see Definition 2.2).
Theorem 3.1.
Suppose there is a protocol that implements $U^{-1}$ to within the average-case distance error $\epsilon$, using $n$ queries to the unknown $d$-dimensional unitary $U$. Then, it must satisfy $n\geq d(d+1)(1-\epsilon)-(d+1)$.
The proof is deferred to Section 3.1. Note that this directly implies a tight lower bound for time-reversal with diamond norm error.
Corollary 3.2.
If we change the average-case distance in Theorem 3.1 to diamond norm, then we have $n\geq d^{2}(1-\epsilon)-1$.
Proof.
By Equation 8, we know that any $\epsilon$-approximate protocol in diamond norm is an $(\frac{d}{d+1}\epsilon)$-approximate protocol in the average-case distance. Therefore, the lower bound in Theorem 3.1 implies a lower bound $d^{2}(1-\epsilon)-1$ for any diamond norm protocol. ∎
From Corollary 3.2, we can further obtain a tight lower bound for the generalized unitary time-reversal task. Specifically, the task is to approximate the time-reverse unitary $U^{-t}=e^{-iHt}$ given a reversing time $t>0$, where $H$ is allowed to be any specific Hamiltonian that satisfies $e^{iH}=U$ and $\left\lVert H\right\rVert\leq\pi$.
Corollary 3.3.
Suppose there is a protocol that implements $U^{-t}$ to within constant diamond norm error for constant $t>0$, using $n$ queries to the unknown $d$-dimensional unitary $U$. Then, it must satisfy $n=\Omega\left\lparen d^{2}\right\rparen$ ).
The proof of Corollary 3.3 is deferred to Section 3.5.
3.1 Main framework
In this subsection, we introduce our main framework and settings for the proof of Theorem 3.1. Let $d\geq 2$ and $n\geq 1$ be arbitrary but fixed integers. We define a sequence of $32n+$ 3 Hilbert spaces $\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n+2}$ such that $\mathcal{H}_{i}\cong\mathbb{C}^{d}$ for $20\leq i\leq 2n+$ 2.
Figure 4: The overall channel from $\mathcal{H}_{0}$ to $\mathcal{H}_{2n+2}$ approximates the identity channel $\mathcal{I}$ when the quantum comb $R$ is a unitary time-reversal protocol using $n$ queries to $U$, i.e., $R\star C_{U}\approx|I\rangle\!\rangle\!\langle\!\langle I|$.
Then, any protocol, which takes $n$ queries to an unknown $d$-dimensional unitary $U$ and produces a $d$-dimensional quantum channel, can be described by an $(n+1)$-comb $R$ on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n+1})$ where the $i$-th query is inserted on $(\mathcal{H}_{2i-1},\mathcal{H}_{2i})$, and the produced quantum channel is defined on $(\mathcal{H}_{0},\mathcal{H}_{2n+1})$ (see the orange region in Figure 4). Furthermore, we make an additional query to $U$ at the end of the protocol $R$, as shown in Figure 4. Then, if $R$ is an approximate time-reversal protocol, the overall channel should approximate the identity channel.
To formalize this, we first define the following $(n+1)$-comb $C_{U}$, which represents the $n+1$ queries to $U$.
Definition 3.4.
For any unitary $U\in\mathbb{U}_{d}$, we define the $(n+1)$-comb $C_{U}$ on $(\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{2n+2})$ as
$C_{U}\coloneqq\underbrace{|U\rangle\!\rangle\!\langle\!\langle U|\otimes\cdots\otimes|U\rangle\!\rangle\!\langle\!\langle U|}_{n+1},$
where the $i$-th tensor factor $|U\rangle\!\rangle\!\langle\!\langle U|$ is the $1$-comb on $(\mathcal{H}_{2i-1},\mathcal{H}_{2i})$ corresponding to the unitary channel $\mathcal{U}$.
Then, we can see the following fact: if $R$ is a time-reversal protocol, the overall channel should still approximate the identity channel when the unknown unitary $U$ is randomly sampled from the Haar measure, i.e., $R\star\mathop{\bf E\/}_{U}[C_{U}]\approx|I\rangle\!\rangle\!\langle\!\langle I|$. Specifically, we have:
Proposition 3.5.
If an $(n+1)$-comb $R$ on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n+1})$ is a time-reversal protocol with the average-case distance error bounded by $\epsilon$, then we have
$\operatorname{tr}\!\left(|I\rangle\!\rangle\!\langle\!\langle I|\cdot\left(R\star\mathop{\bf E\/}_{U}[C_{U}]\right)\right)\geq d^{2}-d(d+1)\epsilon,$
where $|I\rangle\!\rangle\!\langle\!\langle I|$ is the $1$-comb on $(\mathcal{H}_{0},\mathcal{H}_{2n+2})$ corresponding to the identity channel, $\star$ is the link product (see Definition 2.4), $\mathop{\bf E\/}_{U}$ denotes the expectation over a Haar random unitary $U\in\mathbb{U}_{d}$.
Proof.
For any $U\in\mathbb{U}_{d}$, $R$ can produce a quantum channel that approximates $\mathcal{U}^{-1}$ to within the average-case distance error $\epsilon$. Then, we apply the unitary channel $\mathcal{U}$ after the channel produced by $R$ (see Figure 4). Due to the unitary invariance of the average-case distance (see Equation 6), we obtain a channel that approximates the identity channel $\mathcal{I}$ to within the average-case distance error $\epsilon$, i.e.,
$\mathcal{D}_{\textup{avg}}(\mathcal{E}_{R\star C_{U}},\mathcal{I})\leq\epsilon,$
where $\mathcal{E}_{R\star C_{U}}$ denotes the quantum channel corresponding to the $1$-comb (Choi state) $R\star C_{U}$. By Equation 7, this means
$\frac{1}{d^{2}}\operatorname{tr}\!\left(|I\rangle\!\rangle\!\langle\!\langle I|\cdot\left(R\star C_{U}\right)\right)\geq 1-\frac{d+1}{d}\epsilon.$
Then, by taking the Haar expectation over $U\in\mathbb{U}_{d}$ and noting that $\star$ is bilinear, we have
$\operatorname{tr}\!\left(|I\rangle\!\rangle\!\langle\!\langle I|\cdot\left(R\star\mathop{\bf E\/}_{U}[C_{U}]\right)\right)\geq d^{2}-d(d+1)\epsilon.$
On the other hand, we prove the following result, showing that when $U$ is randomly sampled from the Haar measure and the number of queries $n$ is not too large, the overall channel must be very depolarizing (i.e., has low unitarity [CWLY23]). The proof is deferred to Section 3.2.
Theorem 3.6.
For any $(n+1)$-comb $X$ on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2n+1})$, we have
$X\star\mathop{\bf E\/}_{U}[C_{U}]\sqsubseteq\frac{n+1}{d}\cdot I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{0}},$
where $\mathop{\bf E\/}_{U}$ denotes the expectation over a Haar random unitary $U\in\mathbb{U}_{d}$.
Now, we are able to prove our main lower bound (i.e., Theorem 3.1), assuming Theorem 3.6.
Proof of Theorem 3.1.
Combining Proposition 3.5 and Theorem 3.6, we know that
$n+1\geq\operatorname{tr}\!\left(|I\rangle\!\rangle\!\langle\!\langle I|\cdot\left(R\star\mathop{\bf E\/}_{U}[C_{U}]\right)\right)\geq d^{2}-d(d+1)\epsilon,$
which means
$n\geq d(d+1)(1-\epsilon)-(d+1).$
3.2 Proof of Theorem 3.6
Here, we define the following linear operators $A_{k}$ for $1\leq k\leq n$, which we refer to as stair operators and use as a tool in analyzing the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$.
Definition 3.7 (Stair operators).
For each $1\leq k\leq n$, we define the linear operator $A_{k}$ on
$\left(\left(\bigotimes_{i=1}^{k}\mathcal{H}_{2i}\right)\otimes\mathcal{H}_{2n+2}\right)\otimes\left(\bigotimes_{i=1}^{k}\mathcal{H}_{2i-1}\right)\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{k+1}\times\mathfrak{S}_{k}}}{{\cong}}\bigoplus_{\begin{subarray}{c}\lambda\vdash k+1\\ \mu\vdash k\end{subarray}}\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu},$
where $\mathfrak{S}_{k+1}$ acts on $\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\otimes\mathcal{H}_{2k}\otimes\mathcal{H}_{2n+2}$ and $\mathfrak{S}_{k}$ acts on $\mathcal{H}_{1}\otimes\mathcal{H}_{3}\otimes\cdots\otimes\mathcal{H}_{2k-1}$, as
$A_{k}\coloneqq\bigoplus_{\lambda\vdash k+1}\bigoplus_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\lambda,\mu)\end{subarray}}\lvert T\rangle\!\langle S\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert,$
where $T^{\downarrow}$ denotes the standard Young tableau obtained from $T$ by removing the box containing the largest integer, $\textup{Tab}(\lambda,\mu)$ denotes the set of standard Young tableaux $T\in\textup{Tab}(\lambda)$ such that $\textup{Sh}(T^{\downarrow})=\mu$.
Note that each stair operator $A_{k}$ is defined in an asymmetric bipartite system where one of the parties contains $k+1$ copies of $\mathbb{C}^{d}$ and the another contains $k$ copies of $\mathbb{C}^{d}$. In addition, we will demonstrate a contraction process of $A_{k}$ which leads to a smaller $A_{k-1}$. As shown in Figure 5(b), each $A_{k}$ resembles a stair, and the contraction process resembles walking down a stair. This is why we refer to $\{A_{k}\}_{k=1}^{n}$ as stair operators. A similar construction is used in [YKS+24], where they proved a lower bound $\Omega(d)$ for the unitary time-reversal task.
Then, we need the following lemmas about the stair operators, which is used to analyze the behavior of the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$ when combined with another quantum comb. The proofs are deferred to Section 3.3 and Section 3.4. First, the stair operators provide a good upper bound (w.r.t. Löwner order) for $\mathop{\bf E\/}_{U}[C_{U}]$, see also Figure 5(a).
Lemma 3.8.
Let $C_{U}$ be that defined in Definition 3.4. Then
$\mathop{\bf E\/}_{U}[C_{U}]\sqsubseteq I_{\mathcal{H}_{2n+1}}\otimes A_{n},$
where $I_{\mathcal{H}_{2n+1}}$ is the identity operator on $\mathcal{H}_{2n+1}$.
Moreover, the stair operators share the similar tensor contraction property with quantum combs (see Equation 9). Specifically, tracing out the “second-last” subsystem of a stair operator yields an operator proportional to the tensor product of the identity and a reduced stair operator, see also Figure 5(b).
Lemma 3.9.
For any $2\leq k\leq n$, we have
$\operatorname{tr}_{\mathcal{H}_{2k}}(A_{k})=\frac{k+1}{k}\cdot I_{\mathcal{H}_{2k-1}}\otimes A_{k-1},$ (24)
where $I_{\mathcal{H}_{2k-1}}$ is the identity operator on $\mathcal{H}_{2k-1}$. Moreover, we have
$\operatorname{tr}_{\mathcal{H}_{2}}(A_{1})=\frac{2}{d}\cdot I_{\mathcal{H}_{1}}\otimes I_{\mathcal{H}_{2n+2}}.$ (25)
(a) Bounding the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$.
(b) Contraction of the stair operators.
Figure 5: Properties of the stair operators $\{A_{k}\}_{k=1}^{n}$ (see Lemma 3.8 and Lemma 3.9).
Lemma 3.9 implies that when a stair operator is combined with another quantum comb, they can be sequentially contracted. The following corollary is a direct application of Lemma 3.9, which characterizes the behavior of the stair operator $A_{k}$ when combined with an arbitrary $(k+1)$-comb.
Corollary 3.10.
For any $1\leq k\leq n$, and any $(k+1)$-comb $X$ on $(\mathcal{H}_{0},\mathcal{H}_{1},\ldots,\mathcal{H}_{2k+1})$, we have
$X\star(I_{\mathcal{H}_{2k+1}}\otimes A_{k})=\frac{k+1}{d}\cdot I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{0}}.$
Proof.
We use induction on $k$. For the case $k=1$, we have
$X\star(I_{\mathcal{H}_{3}}\otimes A_{1})$ $=\operatorname{tr}_{\mathcal{H}_{1:3}}\!\left(X^{\textup{T}_{\mathcal{H}_{1:3}}}\cdot(I_{\mathcal{H}_{3}}\otimes A_{1})\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2}}\!\left(\operatorname{tr}_{\mathcal{H}_{3}}(X)^{\textup{T}_{\mathcal{H}_{1:2}}}\cdot A_{1}\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2}}\!\left(\left(I_{\mathcal{H}_{2}}\otimes X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1:2}}}\cdot A_{1}\right)$ (26)
$=\operatorname{tr}_{\mathcal{H}_{1}}\left(\left(X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1}}}\cdot\operatorname{tr}_{\mathcal{H}_{2}}(A_{1})\right)$
$=\frac{2}{d}\cdot\operatorname{tr}_{\mathcal{H}_{1}}\left(\left(X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1}}}\cdot I_{\mathcal{H}_{1}}\otimes I_{\mathcal{H}_{2n+2}}\right)$ (27)
$=\frac{2}{d}\cdot I_{\mathcal{H}_{0}}\otimes I_{\mathcal{H}_{2n+2}},$ (28)
where Equation 26 is because $X$ is a $2$-comb and thus $X^{\prime}$ is a $1$-comb, Equation 27 is due to Equation 25 in Lemma 3.9, and Equation 28 is because $X^{\prime}$ is a $1$-comb.
Now, suppose the induction hypothesis holds for $k-1$, we prove it also hods for $k$.
$X\star(I_{\mathcal{H}_{2k+1}}\otimes A_{k})$ $=\operatorname{tr}_{\mathcal{H}_{1:2k+1}}\!\left(X^{\textup{T}_{\mathcal{H}_{1:2k+1}}}\cdot(I_{\mathcal{H}_{2k+1}}\otimes A_{k})\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2k}}\!\left(\operatorname{tr}_{\mathcal{H}_{2k+1}}\!\left(X\right)^{\textup{T}_{\mathcal{H}_{1:2k}}}\cdot A_{k}\right)$
$=\operatorname{tr}_{\mathcal{H}_{1:2k}}\!\left(\left(I_{\mathcal{H}_{2k}}\otimes X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1:2k}}}\cdot A_{k}\right)$ (29)
$=\operatorname{tr}_{\mathcal{H}_{1:2k-1}}\left(\left(X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1:2k-1}}}\cdot\operatorname{tr}_{\mathcal{H}_{2k}}(A_{k})\right)$
$=\frac{k+1}{k}\cdot\operatorname{tr}_{\mathcal{H}_{1:2k-1}}\left(\left(X^{\prime}\right)^{\textup{T}_{\mathcal{H}_{1:2k-1}}}\cdot I_{\mathcal{H}_{2k-1}}\otimes A_{k-1}\right)$ (30)
$=\frac{k+1}{d}\cdot I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{1}},$ (31)
where Equation 29 is because $X$ is an $(k+1)$-comb and thus $X^{\prime}$ is a $k$-comb, Equation 30 is due to Equation 24 in Lemma 3.9, and Equation 31 is due to the induction hypothesis for the case of $k-1$
Then, we are able to prove Theorem 3.6, assuming Lemma 3.8 and Lemma 3.9.
Proof of Theorem 3.6.
Note that
$X\star\mathop{\bf E\/}_{U}[C_{U}]$ $\sqsubseteq X\star(I_{\mathcal{H}_{2n+1}}\otimes A_{n})$ (32)
$=\frac{n+1}{d}\cdot I_{\mathcal{H}_{2n+2}}\otimes I_{\mathcal{H}_{0}},$ (33)
where Equation 32 is by Lemma 3.8 combined with Equation 10, and Equation 33 is by taking $k=n$ in Corollary 3.10. ∎
3.3 Bounding the Haar moment: proof of Lemma 3.8
In this subsection, we bound the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$ using the stair operators (see Lemma 3.8).
First, we need the following fact:
$\mathop{\bf E\/}_{U}[C_{U}]$ $=\bigoplus_{\lambda\vdash n+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|,$
which gives a representation-theoretic expression of the Haar moment $\mathop{\bf E\/}_{U}[C_{U}]$ (see Proposition 4.7 and the proof there). This result is also used in the study of entanglement and bipartite quantum state (see, e.g., [MH07, CWZ24]).
Then, we provide our proof to Lemma 3.8.
Proof of Lemma 3.8.
By Proposition 4.7, we can express $\mathop{\bf E\/}_{U}[C_{U}]$ in the Schur-Weyl basis
$\bigoplus_{\begin{subarray}{c}\lambda\vdash n+1\\ \mu\vdash n+1\end{subarray}}\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu}\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{n+1}\times\mathfrak{S}_{n+1}}}{{\cong}}\bigotimes_{i=1}^{n+1}\mathcal{H}_{2i}\otimes\bigotimes_{i=1}^{n+1}\mathcal{H}_{2i-1}.$
On the other hand, by Definition 3.7, $A_{n}$ is defined on the Schur-Weyl basis
$\bigoplus_{\begin{subarray}{c}\lambda\vdash n+1\\ \mu\vdash n\end{subarray}}\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu}\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{n+1}\times\mathfrak{S}_{n}}}{{\cong}}\bigotimes_{i=1}^{n+1}\mathcal{H}_{2i}\otimes\bigotimes_{i=1}^{n}\mathcal{H}_{2i-1}.$
By comparing these two forms, we note that it is natural to study that for $\mu\vdash n$ and $T,S\in\textup{Tab}(\mu)$, how the linear operator $I_{\mathcal{H}_{2n+1}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle\!\langle S\rvert\in\mathcal{L}(\mathcal{H}_{2n+1}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\mu})\subseteq\mathcal{L}(\bigotimes_{i=1}^{n+1}\mathcal{H}_{2i-1})$ is expressed as an operator on $\bigoplus_{\mu\vdash n+1}\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\mu}$. For this, we have the following result by using Lemma 4.1:
$I_{\mathcal{H}_{2n+1}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle\!\langle S\rvert=\bigoplus_{\nu:\mu\nearrow\nu}I_{\mathcal{Q}_{\nu}}\otimes\lvert T^{\uparrow\nu}\rangle\!\langle S^{\uparrow\nu}\rvert,$
where $T^{\uparrow\nu}$ denotes the standard Young tableau obtained from $T$ by adding a box filled with $n+1$, resulting in the shape $\nu$. This means
$I_{\mathcal{H}_{2n+1}}\otimes A_{n}$ $=\bigoplus_{\lambda\vdash k+1}\bigoplus_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes\bigoplus_{\nu:\mu\nearrow\nu}I_{\mathcal{Q}_{\nu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\lambda,\mu)\end{subarray}}\lvert T\rangle\!\langle S\rvert\otimes\lvert(T^{\downarrow})^{\uparrow\nu}\rangle\!\langle(S^{\downarrow})^{\uparrow\nu}\rvert$
$\sqsupseteq\bigoplus_{\lambda\vdash k+1}\bigoplus_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\lambda,\mu)\end{subarray}}\lvert T\rangle\!\langle S\rvert\otimes\lvert T\rangle\!\langle S\rvert$ (34)
$=\bigoplus_{\lambda\vdash k+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes\sum_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})}\lvert\Phi_{\mu}^{\lambda}\rangle\!\langle\Phi_{\mu}^{\lambda}\rvert$ (35)
$\sqsupseteq\bigoplus_{\lambda\vdash k+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes\left(\sum_{\mu:\mu\nearrow\lambda}\lvert\Phi_{\mu}^{\lambda}\rangle\right)\left(\sum_{\mu:\mu\nearrow\lambda}\langle\Phi_{\mu}^{\lambda}\rvert\right)$ (36)
$=\bigoplus_{\lambda\vdash k+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|$ (37)
$=\mathop{\bf E\/}_{U}[C_{U}],$
where Equation 34 is by discarding those $\nu$ such that $\nu\neq\lambda$; in Equation 35, $\lvert\Phi_{\mu}^{\lambda}\rangle\coloneqq\sum_{\begin{subarray}{c}T\in\textup{Tab}(\lambda,\mu)\end{subarray}}\lvert T\rangle\lvert T\rangle$; Equation 37 is because $|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle=\sum_{T\in\textup{Tab}(\lambda)}\lvert T\rangle\lvert T\rangle=\sum_{\mu:\mu\nearrow\lambda}\lvert\Phi_{\mu}^{\lambda}\rangle$; Equation 36 is because by using Lemma 4.10 and the fact:
$\left(\sum_{\mu:\mu\nearrow\lambda}\langle\Phi_{\mu}^{\lambda}\rvert\right)\left(\sum_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})}\lvert\Phi_{\mu}^{\lambda}\rangle\!\langle\Phi_{\mu}^{\lambda}\rvert\right)^{-1}\left(\sum_{\mu:\mu\nearrow\lambda}\lvert\Phi_{\mu}^{\lambda}\rangle\right)$
$=$ $\left(\sum_{\mu:\mu\nearrow\lambda}\langle\Phi_{\mu}^{\lambda}\rvert\right)\left(\sum_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\mu})}{\dim(\mathcal{P}_{\lambda})}\cdot\frac{1}{\dim(\mathcal{P}_{\mu})^{2}}\cdot\lvert\Phi_{\mu}^{\lambda}\rangle\!\langle\Phi_{\mu}^{\lambda}\rvert\right)\left(\sum_{\mu:\mu\nearrow\lambda}\lvert\Phi_{\mu}^{\lambda}\rangle\right)$ (38)
$=$ $\sum_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\mu})}{\dim(\mathcal{P}_{\lambda})}\cdot\frac{1}{\dim(\mathcal{P}_{\mu})^{2}}\cdot\dim(\mathcal{P}_{\mu})^{2}$
$=$ $1,$ (39)
in which Equation 38 is because $\lvert\Phi_{\mu}^{\lambda}\rangle$ are pairwise orthogonal and $|\textup{Tab}(\lambda,\mu)|=|\textup{Tab}(\mu)|=\dim(\mathcal{P}_{\mu})$, Equation 39 is by the fact that $\sum_{\mu:\mu\nearrow\lambda}\dim(\mathcal{P}_{\mu})=\dim(\mathcal{P}_{\lambda})$. ∎
3.4 Contraction of stair operators: proof of Lemma 3.9
In this subsection, we prove the contraction properties of the stair operators (see Lemma 3.9).
3.4.1 The case of $2\leq k\leq n$
First, we prove for $2\leq k\leq n$,
$\operatorname{tr}_{\mathcal{H}_{2k}}(A_{k})=\frac{k+1}{k}\cdot I_{\mathcal{H}_{2k-1}}\otimes A_{k-1}.$
Note that in the definition of $A_{k}$ (see Definition 3.7), the symmetric group $\mathfrak{S}_{k+1}$ acts on $\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\otimes\mathcal{H}_{2k}\otimes\mathcal{H}_{2n+2}$. We can not directly use Lemma 4.2 since $\operatorname{tr}_{\mathcal{H}_{2k}}$ is tracing out the second-to-last subsystem $\mathcal{H}_{2k}$ instead of the last subsystem $\mathcal{H}_{2n+2}$. For this, we swap the last two tensor factors of $A_{k}$ by using the $k$-th adjacent transposition $s_{k}$. Specifically, let $P(\pi)$ be the action of $\pi\in\mathfrak{S}_{k+1}$ on $\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\otimes\mathcal{H}_{2k}\otimes\mathcal{H}_{2n+2}$. Then, $P(s_{k})A_{k}P(s_{k})$ is the operator on $\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\otimes\mathcal{H}_{2k-2}\otimes\mathcal{H}_{2n+2}\otimes\mathcal{H}_{2k}$ obtained from $A_{k}$ by swapping the positions of $\mathcal{H}_{2k}$ and $\mathcal{H}_{2n+2}$. By the Young’s orthogonal form (see Equation 23), we have
$P(s_{k})A_{k}P(s_{k})$ $=\bigoplus_{\lambda\vdash k+1}\bigoplus_{\mu:\mu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\lambda,\mu)\end{subarray}}\bot_{T,S}\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert,$ (40)
where $\bot_{T,S}$ is defined as
$\frac{1}{r(T)r(S)}\left(\lvert T\rangle\!\langle S\rvert+\sqrt{(r(T)^{2}-1)(r(S)^{2}-1)}\lvert s_{k}T\rangle\!\langle s_{k}S\rvert+\sqrt{r(T)^{2}-1}\lvert s_{k}T\rangle\!\langle S\rvert+\sqrt{r(S)^{2}-1}\lvert T\rangle\!\langle s_{k}S\rvert\right),$ (41)
where $r(T)$ is the axial distance from the $(k+1)$-box to the $k$-box in $T$. Then, we can use Lemma 4.2 to calculate the partial trace $\operatorname{tr}_{\mathcal{H}_{2k}}(P(s_{k})A_{k}P(s_{k}))$.
To this end, let us consider each single summand in the RHS of Equation 40. Suppose $\lambda\vdash k+1$, $\mu\nearrow\lambda$, $T,S\in\textup{Tab}(\lambda,\mu)$. First note that it is not possible that $\textup{Sh}((s_{k}T)^{\downarrow})=\textup{Sh}(S^{\downarrow})$ since we know $\textup{Sh}(T^{\downarrow})=\textup{Sh}(S^{\downarrow})$. Thus by Lemma 4.2, we can ignore the terms $\lvert s_{k}T\rangle\!\langle S\rvert$ and $\lvert T\rangle\!\langle s_{k}S\rvert$ in Equation 41 since they do not contribute after the partial trace. Next, we consider the terms $\lvert s_{k}T\rangle\!\langle s_{k}S\rvert$ and $\lvert T\rangle\!\langle S\rvert$ separately.
Summands involving the term $\lvert s_{k}T\rangle\!\langle s_{k}S\rvert$.
Consider the term $\lvert s_{k}T\rangle\!\langle s_{k}S\rvert$. We have
$\operatorname{tr}_{\mathcal{H}_{2k}}\!\left(\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\frac{\sqrt{(r(T)^{2}-1)(r(S)^{2}-1)}}{r(T)r(S)}\lvert s_{k}T\rangle\!\langle s_{k}S\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\right)$
$\begin{split}=&\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\textup{Sh}((s_{k}T)^{\downarrow})})}I_{\mathcal{Q}_{\textup{Sh}((s_{k}T)^{\downarrow})}}\otimes I_{\mathcal{Q}_{\mu}}\\ &\quad\quad\quad\quad\quad\quad\otimes\mathbbm{1}_{\textup{Sh}((s_{k}T)^{\downarrow})=\textup{Sh}((s_{k}S)^{\downarrow})}\cdot\frac{\sqrt{(r(T)^{2}-1)(r(S)^{2}-1)}}{r(T)r(S)}\lvert(s_{k}T)^{\downarrow}\rangle\!\langle(s_{k}S)^{\downarrow}\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert.\end{split}$ (42)
Then, we consider their summation. For those summands of the form in Equation 42, by re-naming the variables $\textup{Sh}((s_{k}T)^{\downarrow})$ to $\nu$, $(s_{k}T)^{\downarrow}$ to $T$ and $(s_{k}S)^{\downarrow}$ to $S$, we can write the corresponding summation as
$\bigoplus_{\begin{subarray}{c}\nu\vdash k\end{subarray}}\frac{1}{\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes\bigoplus_{\lambda:\nu\nearrow\lambda}\bigoplus_{\begin{subarray}{c}\mu:\mu\nearrow\lambda\\ \mu\neq\nu\end{subarray}}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})}I_{\mathcal{Q}_{\mu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu,\mu\cap\nu)\end{subarray}}\frac{\sqrt{(r(T^{\uparrow\lambda})^{2}-1)(r(S^{\uparrow\lambda})^{2}-1)}}{r(T^{\uparrow\lambda})r(S^{\uparrow\lambda})}\lvert T\rangle\!\langle S\rvert\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert,$ (43)
where $T_{\mu}$ denotes the standard Young tableau obtained from $T$ by moving the largest box of $T$ to the position that results in shape $\mu$ (such moving always exists since $\mu$ is adjacent to $\nu$). Since the summation is taken over all $\mu$ that is adjacent to $\nu$, by re-naming $\mu\cap\nu$ to $\tau$, we can write Equation 43 in an equivalent form:
$\bigoplus_{\begin{subarray}{c}\nu\vdash k\end{subarray}}\frac{1}{\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes\bigoplus_{\tau:\tau\nearrow\nu}\bigoplus_{\begin{subarray}{c}\mu:\tau\nearrow\mu\\ \mu\neq\nu\end{subarray}}\frac{\dim(\mathcal{P}_{\mu\cup\nu})}{\dim(\mathcal{P}_{\mu})}I_{\mathcal{Q}_{\mu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu,\tau)\end{subarray}}\frac{\sqrt{(r(T^{\uparrow\mu\cup\nu})^{2}-1)(r(S^{\uparrow\mu\cup\nu})^{2}-1)}}{r(T^{\uparrow\mu\cup\nu})r(S^{\uparrow\mu\cup\nu})}\lvert T\rangle\!\langle S\rvert\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert.$ (44)
Note that for $T,S\in\textup{Tab}(\nu,\tau)$, $r(T^{\uparrow\mu\cup\nu})=r(S^{\uparrow\mu\cup\nu})=c(\mu\setminus\nu)-c(\nu\setminus\tau)$. Then, by Lemma 4.3, we can write Equation 44 as
$\frac{k+1}{k}\bigoplus_{\begin{subarray}{c}\nu\vdash k\end{subarray}}\frac{1}{\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes\bigoplus_{\tau:\tau\nearrow\nu}\bigoplus_{\begin{subarray}{c}\mu:\tau\nearrow\mu\\ \mu\neq\nu\end{subarray}}\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})}I_{\mathcal{Q}_{\mu}}\otimes\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu,\tau)\end{subarray}}\lvert T\rangle\!\langle S\rvert\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert.$ (45)
Summands involving the term $\lvert T\rangle\!\langle S\rvert$.
Then, consider the term $\lvert T\rangle\!\langle S\rvert$. We have
$\operatorname{tr}_{\mathcal{H}_{2k}}\!\left(\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\frac{1}{r(T)r(S)}\lvert T\rangle\!\langle S\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\right)$
$=$ $\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\mu})\dim(\mathcal{Q}_{\mu})}I_{\mathcal{Q}_{\mu}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\frac{1}{r(T)r(S)}\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert.$ (46)
For those summands of the form in Equation 46, by re-naming the variables $\mu$ to $\nu$, $T^{\downarrow}$ to $T$ and $S^{\downarrow}$ to $S$, we can write the corresponding summation as
$\bigoplus_{\nu\vdash k}\frac{1}{\dim(\mathcal{P}_{\nu})\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\sum_{\begin{subarray}{c}\lambda:\nu\nearrow\lambda\end{subarray}}\,\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu)\end{subarray}}\frac{\dim(\mathcal{P}_{\lambda})}{r(T^{\uparrow\lambda})r(S^{\uparrow\lambda})}\lvert T\rangle\!\langle S\rvert\otimes\lvert T\rangle\!\langle S\rvert.$ (47)
Then, note that $\textup{Tab}(\nu)=\bigcup_{\tau\nearrow\nu}\textup{Tab}(\nu,\tau)$. we can write Equation 47 as
$\bigoplus_{\nu\vdash k}\frac{1}{\dim(\mathcal{P}_{\nu})\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\sum_{\begin{subarray}{c}\lambda:\nu\nearrow\lambda\end{subarray}}\,\sum_{\begin{subarray}{c}\tau:\tau\nearrow\nu\\ \kappa:\kappa\nearrow\nu\end{subarray}}\,\sum_{\begin{subarray}{c}T\in\textup{Tab}(\nu,\tau)\\ S\in\textup{Tab}(\nu,\kappa)\end{subarray}}\frac{\dim(\mathcal{P}_{\lambda})}{r(T^{\uparrow\lambda})r(S^{\uparrow\lambda})}\lvert T\rangle\!\langle S\rvert\otimes\lvert T\rangle\!\langle S\rvert.$ (48)
Note that $r(T^{\uparrow\lambda})=c(\lambda\setminus\nu)-c(\nu\setminus\tau)$ and $r(S^{\uparrow\lambda})=c(\lambda\setminus\nu)-c(\nu\setminus\kappa)$. By Lemma 4.4, those terms such that $\tau\neq\kappa$ can be discarded. Therefore, Equation 48 can be simplified to:
$\bigoplus_{\nu\vdash k}\frac{1}{\dim(\mathcal{P}_{\nu})\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\sum_{\begin{subarray}{c}\lambda:\nu\nearrow\lambda\end{subarray}}\,\sum_{\tau:\tau\nearrow\nu}\,\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu,\tau)\end{subarray}}\frac{\dim(\mathcal{P}_{\lambda})}{r(T^{\uparrow\lambda})r(S^{\uparrow\lambda})}\lvert T\rangle\!\langle S\rvert\otimes\lvert T\rangle\!\langle S\rvert.$ (49)
Then, by Lemma 4.5, we can further simplify Equation 49 to
$\frac{k+1}{k}\cdot\bigoplus_{\nu\vdash k}\frac{1}{\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\sum_{\tau:\tau\nearrow\nu}\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})}\sum_{\begin{subarray}{c}T,S\in\textup{Tab}(\nu,\tau)\end{subarray}}\lvert T\rangle\!\langle S\rvert\otimes\lvert T\rangle\!\langle S\rvert.$ (50)
Putting it all together.
Then, combining the above results, we can see that the sum of Equation 45 and Equation 50 is:
$\frac{k+1}{k}\cdot\bigoplus_{\nu\vdash k}\frac{1}{\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes\bigoplus_{\tau:\tau\nearrow\nu}\bigoplus_{\mu:\tau\nearrow\mu}\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})}I_{\mathcal{Q}_{\mu}}\otimes\sum_{T,S\in\textup{Tab}(\nu,\tau)}\lvert T\rangle\!\langle S\rvert\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert.$ (51)
Note that in Equation 51, $\bigoplus_{\mu:\tau\nearrow\mu}I_{\mathcal{Q}_{\mu}}\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert$ is an operator on $\bigotimes_{i=1}^{k}\mathcal{H}_{2i-1}$. Thus Lemma 4.1 implies that
$\bigoplus_{\mu:\tau\nearrow\mu}I_{\mathcal{Q}_{\mu}}\otimes\lvert T_{\mu}\rangle\!\langle S_{\mu}\rvert=I_{\mathcal{H}_{2k-1}}\otimes I_{\mathcal{Q}_{\tau}}\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert.$
This means Equation 51 is
$\frac{k+1}{k}I_{\mathcal{H}_{2k-1}}\otimes\bigoplus_{\nu\vdash k}\bigoplus_{\tau:\tau\nearrow\nu}\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})\dim(\mathcal{Q}_{\nu})}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\tau}}\otimes\sum_{T,S\in\textup{Tab}(\nu,\tau)}\lvert T\rangle\!\langle S\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert$
$=$ $\frac{k+1}{k}\cdot I_{\mathcal{H}_{2k-1}}\otimes A_{k-1}.$
3.4.2 The case of $k=1$
Then, we prove
$\operatorname{tr}_{\mathcal{H}_{2}}(A_{1})=\frac{2}{d}\cdot I_{\mathcal{H}_{1}}\otimes I_{\mathcal{H}_{2n+2}}.$
by explicit calculation. Specifically, let λ=
{ytableau}
[…]
{ytableau}
[…]
and μ=
{ytableau}
[…]
{ytableau}
[…]
{ytableau}
1&2
[…]
{ytableau}
1&2
[…]
and S=
{ytableau}
1
2
[…]
{ytableau}
1
2
[…]
. Thus, $\nu\nearrow\lambda$, $\nu\nearrow\mu$, $\dim(\mathcal{P}_{\lambda})=\dim(\mathcal{P}_{\mu})=\dim(\mathcal{P}_{\nu})=1$, $\dim(\mathcal{Q}_{\lambda})=\frac{d(d+1)}{2}$, $\dim(\mathcal{Q}_{\mu})=\frac{d(d-1)}{2}$ and $\dim(\mathcal{Q}_{\nu})=d$. Then, we have
$A_{1}$ $=\frac{2}{d(d+1)}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\lvert T\rangle\!\langle T\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert+\frac{2}{d(d-1)}I_{\mathcal{Q}_{\mu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\lvert S\rangle\!\langle S\rvert\otimes\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert.$
Next, we swap the subsystems $\mathcal{H}_{2}$ and $\mathcal{H}_{2n+2}$. In fact, note that the swap $s_{1}$ acts trivially on $\lvert T\rangle$ and $\lvert S\rangle$. Thus, we can directly calculate the partial trace on $A_{1}$ by Lemma 4.2:
$\operatorname{tr}_{\mathcal{H}_{2}}(A_{1})$ $=\frac{1}{d}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert\otimes\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert+\frac{1}{d}I_{\mathcal{Q}_{\nu}}\otimes I_{\mathcal{Q}_{\nu}}\otimes\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\otimes\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert$
$=\frac{2}{d}I_{\mathcal{H}_{1}}\otimes I_{\mathcal{H}_{2n+2}},$ (52)
where Equation 52 is because T=S=
{ytableau}
1
[…]
{ytableau}
1
[…]
and thus $\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert=\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert=I_{\mathcal{P}_{\nu}}$.
3.5 Generalized time-reversal
In this section, let us prove Corollary 3.3, which gives a lower bound for generalized time-reversal of an unknown unitary.
Proof of Corollary 3.3.
Assume for contradiction that there is a protocol $\mathcal{A}$ to implement a unitary $U^{-t}=e^{-iHt}$ to within diamond norm error $\epsilon=O(1)$ for some $t=O(1)$ and $H$ that satisfies $e^{iH}=U$ and $\left\lVert H\right\rVert\leq\pi$, using $o\left\lparen d^{2}\right\rparen$ ) queries to $U$. Let us choose $b=\lceil\log(1/\epsilon)\rceil$. Then, it is easy to find $t^{\prime}=k+\frac{a}{b}$ for some $a<b\in\mathbb{N}$ and $k\in\mathbb{N}$, such that $\left\lvert t^{\prime}-t\right\rvert\leq\epsilon$. Note that we can bound
$\left\lVert e^{-iHt}-e^{-iHt^{\prime}}\right\rVert_{\diamond}\leq 2\pi\epsilon.$ (53)
To see Equation 53, we first use the inequality between the diamond norm and the operator norm
$\left\lVert e^{-iHt}-e^{-iHt^{\prime}}\right\rVert_{\diamond}\leq 2\left\lVert e^{-iHt}-e^{-iHt^{\prime}}\right\rVert.$ (54)
To further bound the operator norm on the RHS of Equation 54, we observe
$\left\lVert e^{-iHt}-e^{-iHt^{\prime}}\right\rVert=\left\lVert e^{-iHt}\left\lparen I-e^{-iH\left\lvert t-t^{\prime}\right\rvert}\right\rparen\right\rVert=\left\lVert I-e^{-iH\left\lvert t-t^{\prime}\right\rvert}\right\rVert,$
which can be upper bounded by $\left\lvert t-t^{\prime}\right\rvert\cdot\left\lVert H\right\rVert$, through looking at the eigenvalues and using $\left\lvert 1-e^{ix}\right\rvert\leq\left\lvert x\right\rvert$. Equation 53 implies that $\mathcal{A}$ is also a protocol to implement $e^{-iHt^{\prime}}$ to within diamond norm error $(2\pi+1)\epsilon$.
Now let us construct to a protocol to approximately implement $U^{-1}$, based on protocol $\mathcal{A}$. First, using $k$ more queries to $U$ followed by $\mathcal{A}$, we can obtain a protocol $\mathcal{B}$ to implement the unitary $e^{-iHa/b}$ to within diamond norm error $(2\pi+1)\epsilon$. Further, if we repeat protocol $\mathcal{B}$ $b$ times, we have a protocol $\mathcal{C}$ to implement the unitary $e^{-iHa}$ to within diamond norm error $b(2\pi+1)\epsilon$. Finally, if we use $a-1$ more queries to $U$ followed by protocol $\mathcal{C}$, we obtain a protocol $\mathcal{D}$ to implement the time-reverse $U^{-1}$ to within diamond norm error $b(2\pi+1)\epsilon$. Since $t=O(1)$ and $\epsilon=O(1)$, we have $k$, $a$ and $b$ also being constant. Consequently, protocol $\mathcal{D}$ only uses $o\left\lparen d^{2}\right\rparen$ ) queries to $U$. This leads to a contradiction with Corollary 3.2, and the conclusion immediately follows. ∎
4 Deferred technical lemmas
4.1 Raising and lowering of Young diagrams
In this subsection, we present two representation-theoretic results concerning the raising and lowering of Young diagrams. These results follow from standard techniques involving the Clebsch-Gordan transform and Schur transform [BCH07, BCH06, Har05] (see also [YSM23, SMKH22]). However, for the reader’s convenience, we provide simple and self-contained proofs here.
Lemma 4.1.
Let $n\geq 2$ be an integer and $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{n}$ be a sequence of Hilbert spaces such that $\mathcal{H}_{i}\cong\mathbb{C}^{d}$ for $1\leq i\leq n$. Let $\mathfrak{S}_{n}$ act on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$ and consider the corresponding decomposition $\bigotimes_{i=1}^{n}\mathcal{H}_{i}\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathfrak{S}_{n}}}{{\cong}}\bigoplus_{\lambda\vdash n}\mathcal{Q}_{\lambda}\otimes\mathcal{P}_{\lambda}$. Then, for $\mu\vdash n-1$, $T,S\in\textup{Tab}(\mu)$, we have
$I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle\!\langle S\rvert=\bigoplus_{\lambda:\mu\nearrow\lambda}I_{\mathcal{Q}_{\lambda}}\otimes\lvert T^{\uparrow\lambda}\rangle\!\langle S^{\uparrow\lambda}\rvert,$
where $T^{\uparrow\lambda}$ denotes the standard Young tableau obtained from $T$ by adding a box filled with $n$, resulting in the shape $\lambda$.
Proof.
We consider the linear operator $I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle\in\mathcal{L}(\mathcal{Q}_{\mu},\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\mu})\subseteq\mathcal{L}(\mathcal{Q}_{\mu},\bigotimes_{i=1}^{n-1}\mathcal{H}_{i})$. By Pieri’s rule:
$\mathbb{C}^{d}\otimes\mathcal{Q}_{\mu}\stackrel{{\scriptstyle\mathbb{U}_{d}}}{{\cong}}\bigoplus_{\lambda:\mu\nearrow\lambda}\mathcal{Q}_{\lambda},$
we can see that
$I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle=\bigoplus_{\lambda:\mu\nearrow\lambda}I_{\mathcal{Q}_{\lambda}}\otimes\lvert T_{\lambda}\rangle,$ (55)
where $\lvert T_{\lambda}\rangle$ is a unit vector and by Schur-Weyl duality on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$, we know that $\lvert T_{\lambda}\rangle\in\mathcal{P}_{\lambda}$. Suppose $T=\mu^{(1)}\rightarrow\cdots\rightarrow\mu^{(n-1)}$ where $\mu^{(n-1)}=\mu$. For an integer $11\leq k\leq n-$ 1, we apply $P(e_{\mu^{(k)}})$ (i.e., the action of $e_{\mu^{(k)}}$ on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$, where $e_{\mu^{(k)}}$ is defined in Equation 20) to both sides of Equation 55. For the LHS, since $e_{\mu^{(k)}}\in\mathbb{C}\mathfrak{S}_{k}\subseteq\mathbb{C}\mathfrak{S}_{n-1}$, $P(e_{\mu^{(k)}})$ acts non-trivially only on $\bigotimes_{i=1}^{n-1}\mathcal{H}_{i}$, thus we have
$P(e_{\mu^{(k)}})(I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle)=I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes e_{\mu^{(k)}}\lvert T\rangle=I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\lvert T\rangle.$
For the RHS, we have
$P(e_{\mu^{(k)}})\left(\bigoplus_{\lambda:\mu\nearrow\lambda}I_{\mathcal{Q}_{\lambda}}\otimes\lvert T_{\lambda}\rangle\right)=\bigoplus_{\lambda:\mu\nearrow\lambda}I_{\mathcal{Q}_{\lambda}}\otimes e_{\mu^{(k)}}\lvert T_{\lambda}\rangle.$
Therefore, we can see that
$e_{\mu^{(k)}}\lvert T_{\lambda}\rangle=\lvert T_{\lambda}\rangle,$
for any $11\leq k\leq n-$ 1. This means $\lvert T_{\lambda}\rangle\in\mathcal{P}_{\lambda}$, when restricted as a vector in a representation of $\mathfrak{S}_{k}$, is a vector in $\mathcal{P}_{\mu^{(k)}}$. Therefore, $\lvert T_{\lambda}\rangle=\lvert\mu^{(1)}\rightarrow\cdots\rightarrow\mu^{(n-1)}\rightarrow\lambda\rangle=\lvert T^{\uparrow\lambda}\rangle$.
We can analogously prove $I_{\mathcal{H}_{n}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\langle S\rvert=\bigoplus_{\lambda:\mu\nearrow\lambda}I_{\mathcal{Q}_{\lambda}}\otimes\langle S^{\uparrow\lambda}\rvert$. This completes the proof. ∎
The following result can be viewed, in some sense, as a “dual” of Lemma 4.1.
Lemma 4.2.
Let $n\geq 2$ be an integer and $\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{n}$ be a sequence of Hilbert spaces such that $\mathcal{H}_{i}\cong\mathbb{C}^{d}$ for $1\leq i\leq n$. Let $\mathfrak{S}_{n}$ act on $\bigotimes_{i=1}^{n}\mathcal{H}_{i}$ and consider the corresponding decomposition $\bigotimes_{i=1}^{n}\mathcal{H}_{i}\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathfrak{S}_{n}}}{{\cong}}\bigoplus_{\lambda\vdash k}\mathcal{Q}_{\lambda}\otimes\mathcal{P}_{\lambda}$. Then, for $\lambda\vdash n$, $T,S\in\textup{Tab}(\lambda)$, we have
$\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert)=\mathbbm{1}_{\textup{Sh}(T^{\downarrow})=\textup{Sh}(S^{\downarrow})}\cdot\frac{\dim(\mathcal{Q}_{\lambda})}{\dim(\mathcal{Q}_{\textup{Sh}(T^{\downarrow})})}\cdot I_{\mathcal{Q}_{\textup{Sh}(T^{\downarrow})}}\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert,$
where $T^{\downarrow}$ denotes the standard Young tableau obtained from $T$ by removing the box containing the largest integer.
Proof.
Let
$W\coloneqq\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert).$ (56)
Note that $W\in\mathcal{L}(\bigotimes_{i=1}^{n-1}\mathcal{H}_{i})$. For any $U\in\mathbb{U}_{d}$, we have
$U^{\otimes n-1}WU^{{\dagger}\otimes n-1}$ $=\operatorname{tr}_{\mathcal{H}_{n}}\!\left(U^{\otimes n}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert)U^{{\dagger}\otimes n}\right)$
$=\operatorname{tr}_{\mathcal{H}_{n}}\!\left(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert\right)$
$=\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert).$
By Schur-Weyl duality on $\bigotimes_{i=1}^{n-1}\mathcal{H}_{i}$ and Schur’s lemma, this means
$W=\bigoplus_{\mu\vdash n-1}I_{\mathcal{Q}_{\mu}}\otimes X_{\mu},$ (57)
for some $X_{\mu}\in\mathcal{L}(\mathcal{P}_{\mu})$. On the other hand, suppose $T=\lambda_{T}^{(1)}\rightarrow\lambda_{T}^{(2)}\rightarrow\cdots\rightarrow\lambda_{T}^{(n)}$ and $S=\lambda_{S}^{(1)}\rightarrow\lambda_{S}^{(2)}\rightarrow\cdots\rightarrow\lambda_{S}^{(n)}$, where $\lambda_{T}^{(n)}=\lambda_{S}^{(n)}=\lambda$. For $k\leq n-1$ and $\nu,\tau\vdash k$, we left-apply $P(e_{\nu})$ (i.e., the action of $e_{\nu}$ on $\bigotimes_{i=1}^{n-1}\mathcal{H}_{i}$, where $e_{\nu}$ is defined in Equation 20) and right-apply $P(e_{\tau})$ to $W$. First, by the definition of $W$ (c.f. Equation 56), we have
$P(e_{\nu})WP(e_{\tau})$ $=\operatorname{tr}_{\mathcal{H}_{n}}\!\left(P(e_{\nu})(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert)P(e_{\tau})\right)$ (58)
$=\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes e_{\nu}\lvert T\rangle\!\langle S\rvert e_{\tau})$
$=\mathbbm{1}_{\nu=\lambda_{T}^{(k)}}\cdot\mathbbm{1}_{\tau=\lambda_{S}^{(k)}}\cdot\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert T\rangle\!\langle S\rvert)$
$=\mathbbm{1}_{\nu=\lambda_{T}^{(k)}}\cdot\mathbbm{1}_{\tau=\lambda_{S}^{(k)}}\cdot W,$ (59)
where Equation 58 is because $e_{\nu},e_{\tau}\in\mathbb{C}\mathfrak{S}_{k}\subseteq\mathbb{C}\mathfrak{S}_{n-1}$ acts non-trivially only on $\bigotimes_{i=1}^{n-1}\mathcal{H}_{i}$. Then, by recursively using Equation 59, we know that
$W=P\!\left(e_{\lambda_{T}^{(1)}}\right)\cdots P\!\left(e_{\lambda_{T}^{(n-1)}}\right)\cdot W\cdot P\!\left(e_{\lambda_{S}^{(n-1)}}\right)\cdots P\!\left(e_{\lambda_{S}^{(1)}}\right).$
On the other hand, by using Equation 57, we have
$W$ $=P\!\left(e_{\lambda_{T}^{(1)}}\right)\cdots P\!\left(e_{\lambda_{T}^{(n-1)}}\right)\cdot W\cdot P\!\left(e_{\lambda_{S}^{(n-1)}}\right)\cdots P\!\left(e_{\lambda_{S}^{(1)}}\right)$
$=\bigoplus_{\mu\vdash n-1}I_{\mathcal{Q}_{\mu}}\otimes\left(e_{\lambda_{T}^{(1)}}\cdots e_{\lambda_{T}^{(n-1)}}\cdot X_{\mu}\cdot e_{\lambda_{S}^{(n-1)}}\cdots e_{\lambda_{S}^{(1)}}\right)$
$=\bigoplus_{\mu\vdash n-1}I_{\mathcal{Q}_{\mu}}\otimes\left(\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert\cdot X_{\mu}\cdot\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\right),$ (60)
where Equation 60 is by using Equation 22. Note that $\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert$ is a linear projector on $\mathcal{P}_{\textup{Sh}(T^{\downarrow})}$ and $\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert$ is a linear projector on $\mathcal{P}_{\textup{Sh}(S^{\downarrow})}$. This means $\lvert T^{\downarrow}\rangle\!\langle T^{\downarrow}\rvert\cdot X_{\mu}\cdot\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert$ is non-zero only when $\mu=\textup{Sh}(T^{\downarrow})=\textup{Sh}(S^{\downarrow})$. Therefore, we have
$W=\mathbbm{1}_{\textup{Sh}(T^{\downarrow})=\textup{Sh}(S^{\downarrow})}\cdot c\cdot I_{\mathcal{Q}_{\textup{Sh}(T^{\downarrow})}}\otimes\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert,$ (61)
for some number $c$.
Now suppose $\textup{Sh}(T^{\downarrow})=\textup{Sh}(S^{\downarrow})$. Thus both $\lvert T^{\downarrow}\rangle$ and $\lvert S^{\downarrow}\rangle$ are in the same irreducible representation $\mathcal{P}_{\textup{Sh}(T^{\downarrow})}$. Let $K\in\mathbb{C}\mathfrak{S}_{n-1}$ such that $K\lvert T^{\downarrow}\rangle=\lvert S^{\downarrow}\rangle$ (such $K$ always exists by the Jacobson density theorem [EGH+11]). Then, $K$ also satisfies $K\lvert T\rangle=\lvert S\rangle$. To see this, we view $\lvert T\rangle$ and $\lvert S\rangle$ as vectors in the restricted representation $\textup{Res}^{\mathfrak{S}_{n}}_{\mathfrak{S}_{n-1}}\mathcal{P}_{\lambda}$. By the definition of Young basis, they both are in the same irreducible representation $\mathcal{P}_{\textup{Sh}(T^{\downarrow})}$ and correspond to $\lvert T^{\downarrow}\rangle$ and $\lvert S^{\downarrow}\rangle$, respectively. Therefore, by the definition of $K$, $K$ maps $\lvert T\rangle$ to $\lvert S\rangle$. Then, consider the trace $\operatorname{tr}(P(K)\cdot W)$, where $P(K)$ is the action of $K$ on $\bigotimes_{i=1}^{n-1}\mathcal{H}_{i}$. On the one hand, by Equation 61, we have
$\operatorname{tr}(P(K)\cdot W)$ $=\operatorname{tr}\!\left(c\cdot I_{\mathcal{Q}_{\textup{Sh}(T^{\downarrow})}}\otimes K\lvert T^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\right)=\operatorname{tr}\!\left(c\cdot I_{\mathcal{Q}_{\textup{Sh}(T^{\downarrow})}}\otimes\lvert S^{\downarrow}\rangle\!\langle S^{\downarrow}\rvert\right)=c\cdot\dim(\mathcal{Q}_{\textup{Sh}(T^{\downarrow})}).$
On the other hand, by Equation 56, we have
$\operatorname{tr}(P(K)\cdot W)$ $=\operatorname{tr}\!\left(\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes K\lvert T\rangle\!\langle S\rvert)\right)=\operatorname{tr}\!\left(\operatorname{tr}_{\mathcal{H}_{n}}(I_{\mathcal{Q}_{\lambda}}\otimes\lvert S\rangle\!\langle S\rvert)\right)=\dim(\mathcal{Q}_{\lambda}).$
Therefore, $c=\dim(\mathcal{Q}_{\lambda})/\dim(\mathcal{Q}_{\textup{Sh}(T^{\downarrow})})$, completing the proof upon substitution into Equation 61. ∎
4.2 Combinatorics on Young diagrams
Here, we provide some combinatorial results on Young diagrams. Our proofs for Lemma 4.4 and Lemma 4.5 are based on Kerov’s interlacing sequences [Ker93].
Lemma 4.3.
Suppose $\mu,\nu\vdash n$, $\mu\neq\nu$ and $\mu,\nu$ are adjacent. Then
$\frac{\dim(\mathcal{P}_{\mu})\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\mu\cup\nu})\dim(\mathcal{P}_{\mu\cap\nu})}=\frac{n}{n+1}\left(1-\frac{1}{(c(\mu\setminus\nu)-c(\nu\setminus\mu))^{2}}\right),$
where $\mu\setminus\nu$ contains only one box and $c(\mu\setminus\nu)$ is the axial coordinate of this box (and similarly for $\nu\setminus\mu$).
Proof.
Let $\square_{\mu}$ be the box in $\mu\setminus\nu$ and $\square_{\nu}$ be the box in $\nu\setminus\mu$. Define $\tau\coloneqq\mu\cap\nu$. Denote by $H_{\lambda}(\square)$ the set of boxes in $\lambda$ that are directly to the left and above the box $\square$, including $\square$ itself (i.e., a hook in the reverse direction). By the hook length formula (see Equation 14), we can easily see that
$\frac{\dim(\mathcal{P}_{\mu})}{\dim(\mathcal{P}_{\tau})}=n\cdot\prod_{\square\in H_{\tau}(\square_{\mu})}\frac{h_{\tau}(\square)}{h_{\tau}(\square)+1}.$
Similarly
$\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})}=n\cdot\prod_{\square\in H_{\tau}(\square_{\nu})}\frac{h_{\tau}(\square)}{h_{\tau}(\square)+1}.$
Then, note that $\square_{\mu},\square_{\nu}$ must be at different rows and columns. Thus, there is exactly one box in $H_{\tau}(\square_{\mu})\cap H_{\tau}(\square_{\nu})$, which we denote by $\widetilde{\square}$. Then, we have
$\frac{\dim(\mathcal{P}_{\mu\cup\nu})}{\dim(\mathcal{P}_{\tau})}=n(n+1)\cdot\left(\prod_{\begin{subarray}{c}\square\in H_{\tau}(\square_{\mu})\\ \square\neq\widetilde{\square}\end{subarray}}\frac{h_{\tau}(\square)}{h_{\tau}(\square)+1}\right)\cdot\left(\prod_{\begin{subarray}{c}\square\in H_{\tau}(\square_{\nu})\\ \square\neq\widetilde{\square}\end{subarray}}\frac{h_{\tau}(\square)}{h_{\tau}(\square)+1}\right)\cdot\left(\frac{h_{\tau}(\widetilde{\square})}{h_{\tau}(\widetilde{\square})+2}\right).$
Then, we have
$\frac{\dim(\mathcal{P}_{\mu})\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\mu\cup\nu})\dim(\mathcal{P}_{\tau})}$ $=\frac{n}{n+1}\cdot\left(\frac{h_{\tau}(\widetilde{\square})}{h_{\tau}(\widetilde{\square})+1}\right)^{2}\cdot\frac{h_{\tau}(\widetilde{\square})+2}{h_{\tau}(\widetilde{\square})}$
$=\frac{n}{n+1}\cdot\left(1-\frac{1}{(h_{\tau}(\widetilde{\square})+1)^{2}}\right)$
$=\frac{n}{n+1}\left(1-\frac{1}{(c(\square_{\mu})-c(\square_{\nu}))^{2}}\right),$ (62)
where Equation 62 is because $h_{\tau}(\widetilde{\square})+1=|c(\square_{\mu})-c(\square_{\nu})|$ (see, e.g., Figure 6(b)).
(a) The sets $H_{\tau}(\square_{\mu})$ and $H_{\tau}(\square_{\nu})$.
(b) The axial distance between $\square_{\mu}$ and $\square_{\nu}$.
Figure 6: An example illustrating the idea in the proof of Lemma 4.3.
Lemma 4.4.
Suppose $\nu\vdash n$, $\tau\nearrow\nu$, $\kappa\nearrow\nu$, $\tau\neq\kappa$. Then,
$\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\tau)}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\kappa)}=0,$
where $\lambda\setminus\nu$ contains only one box and $c(\lambda\setminus\nu)$ is the axial coordinate of this box (and similarly for $\nu\setminus\tau$ and $\nu\setminus\kappa$).
Proof.
Note that
$\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\tau)}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\kappa)}$
$=$ $\frac{1}{c(\nu\setminus\tau)-c(\nu\setminus\kappa)}\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\left(\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\tau)}-\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\kappa)}\right)$
Therefore, it suffices to prove that for any $\tau\nearrow\nu$, we have
$\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\tau)}=0.$
Let $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{L})$ and $\bm{\beta}=(\beta_{1},\ldots,\beta_{L-1})$ be the interlacing sequences of $\nu$. The box in $\nu\setminus\tau$ is a removable box of $\nu$. Thus this box is at $\beta_{m}$ for some $m$ such that $\beta_{m}=c(\nu\setminus\tau)$. Any $\lambda$ such that $\nu\nearrow\lambda$ is obtained by adding a box at an addable position of $\nu$, which corresponds to an $\alpha_{i}$ for $1\leq i\leq L$. Furthermore, for the $\lambda$ obtained by adding a box at $\alpha_{i}$, we know $c(\lambda\setminus\nu)=\alpha_{i}$. Then, by Equation 15, we have
$\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\frac{1}{c(\lambda\setminus\nu)-c(\nu\setminus\tau)}=(n+1)\cdot\sum_{i=1}^{L}\frac{1}{\alpha_{i}-\beta_{m}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}},$
which is 00 by Lemma 4.6. ∎
Lemma 4.5.
Suppose $\nu\vdash n$, $\tau\nearrow\nu$. Then
$\sum_{\lambda:\nu\nearrow\lambda}\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}\frac{1}{(c(\lambda\setminus\nu)-c(\nu\setminus\tau))^{2}}=\frac{n+1}{n}\cdot\frac{\dim(\mathcal{P}_{\nu})}{\dim(\mathcal{P}_{\tau})},$ (63)
where $\lambda\setminus\nu$ contains only one box and $c(\lambda\setminus\nu)$ is the axial coordinate of this box (and similarly for $\nu\setminus\tau$).
Proof.
Let $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{L})$ and $\bm{\beta}=(\beta_{1},\ldots,\beta_{L-1})$ be the interlacing sequences of $\nu$. The box in $\nu\setminus\tau$ is a removable box of $\nu$. Thus this box is at $\beta_{m}$ for some $m$ such that $\beta_{m}=c(\nu\setminus\tau)$. Any $\lambda$ such that $\nu\nearrow\lambda$ is obtained by adding a box at an addable position of $\nu$, which corresponds to an $\alpha_{i}$ for $1\leq i\leq L$. Furthermore, for the $\lambda$ obtained by adding a box at $\alpha_{i}$, we know $c(\lambda\setminus\nu)=\alpha_{i}$. Then, by Equation 15 (applied to $\frac{\dim(\mathcal{P}_{\lambda})}{\dim(\mathcal{P}_{\nu})}$) and Equation 16 (applied to $\frac{\dim(\mathcal{P}_{\tau})}{\dim(\mathcal{P}_{\nu})}$), Equation 63 is equivalent to
$\sum_{i=1}^{L}\frac{1}{(\alpha_{i}-\beta_{m})^{2}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}=\frac{1}{(\alpha_{L}-\beta_{m})(\beta_{m}-\alpha_{1})}\prod_{j=1}^{m-1}\frac{\beta_{m}-\beta_{j}}{\beta_{m}-\alpha_{j+1}}\prod_{j=m+1}^{L-1}\frac{\beta_{m}-\beta_{j}}{\beta_{m}-\alpha_{j}}.$ (64)
Then, we prove Equation 64 by induction on $L$. For the case $L=2$, $m$ must be $1$, and the holding of Equation 64 can be checked by direct calculation. Now, suppose Equation 64 holds for $L$. We want to prove it also holds for $L+1$, i.e., the following holds:
$\sum_{i=1}^{L+1}\frac{1}{(\alpha_{i}-\beta_{m})^{2}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L+1}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}=\frac{1}{(\alpha_{L+1}-\beta_{m})(\beta_{m}-\alpha_{1})}\prod_{j=1}^{m-1}\frac{\beta_{m}-\beta_{j}}{\beta_{m}-\alpha_{j+1}}\prod_{j=m+1}^{L}\frac{\beta_{m}-\beta_{j}}{\beta_{m}-\alpha_{j}}.$ (65)
For the case of $L+1$, we can assume without loss of generality that $m\leq L-1$. This is because if $m=L$, then we consider the reversed interlacing sequences $\bm{\alpha}^{\prime}=(\alpha^{\prime}_{1},\ldots,\alpha^{\prime}_{L+1})$, $\bm{\beta}^{\prime}=(\beta^{\prime}_{1},\ldots,\beta^{\prime}_{L})$ such that $\alpha^{\prime}_{i}=-\alpha_{L+2-i}$ and $\beta^{\prime}_{i}=-\beta_{L+1-i}$. Then, it is easy to see that the holding of Equation 65 on $\bm{\alpha},\bm{\beta}$ with $m=L$ is equivalent to that on $\bm{\alpha}^{\prime}$, $\bm{\beta}^{\prime}$ with $m=1$.
It is easy to see that
[…]
By the induction hypothesis, it suffices to prove
[…]
Then, note that the LHS of Equation 65 can be written as
$\sum_{i=1}^{L}\frac{\alpha_{i}-\beta_{L}}{(\alpha_{i}-\beta_{m})^{2}\cdot(\alpha_{i}-\alpha_{L+1})}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}+\frac{1}{(\alpha_{L+1}-\beta_{m})^{2}}\prod_{j=1}^{L}\frac{\alpha_{L+1}-\beta_{j}}{\alpha_{L+1}-\alpha_{j}}.$
Therefore, the LHS of Equation 66 is:
$\sum_{i=1}^{L}\frac{\beta_{L}-\alpha_{L+1}}{(\alpha_{i}-\beta_{m})(\beta_{m}-\alpha_{L+1})(\alpha_{i}-\alpha_{L+1})}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}+\frac{1}{(\alpha_{L+1}-\beta_{m})^{2}}\prod_{j=1}^{L}\frac{\alpha_{L+1}-\beta_{j}}{\alpha_{L+1}-\alpha_{j}}$
$=$ $\frac{\beta_{L}-\alpha_{L+1}}{(\beta_{m}-\alpha_{L+1})^{2}}\Bigg{(}\sum_{i=1}^{L}\left(\frac{1}{\alpha_{i}-\beta_{m}}-\frac{1}{\alpha_{i}-\alpha_{L+1}}\right)\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}-\frac{1}{\alpha_{L+1}-\alpha_{L}}\prod_{j=1}^{L-1}\frac{\alpha_{L+1}-\beta_{j}}{\alpha_{L+1}-\alpha_{j}}\Bigg{)}$
$=$ $-\frac{\beta_{L}-\alpha_{L+1}}{(\beta_{m}-\alpha_{L+1})^{2}}\Bigg{(}\sum_{i=1}^{L}\frac{1}{\alpha_{i}-\alpha_{L+1}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}+\frac{1}{\alpha_{L+1}-\alpha_{L}}\prod_{j=1}^{L-1}\frac{\alpha_{L+1}-\beta_{j}}{\alpha_{L+1}-\alpha_{j}}\Bigg{)}$ (67)
$=$ $-\frac{\beta_{L}-\alpha_{L+1}}{(\beta_{m}-\alpha_{L+1})^{2}}\Bigg{(}\sum_{i=1}^{L+1}\frac{1}{\alpha_{i}-\beta_{L}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L+1}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}\Bigg{)}$
[…]
where Equation 67 is by Lemma 4.6 and Equation 68 is also by Lemma 4.6. ∎
Lemma 4.6.
Suppose $\alpha_{1},\ldots,\alpha_{L},\beta_{1},\ldots,\beta_{L-1}$ are pairwise distinct. Let $11\leq m\leq L-$ 1. Then we have
$\sum_{i=1}^{L}\frac{1}{\alpha_{i}-\beta_{m}}\prod_{j=1}^{i-1}\frac{\alpha_{i}-\beta_{j}}{\alpha_{i}-\alpha_{j}}\prod_{j=i+1}^{L}\frac{\alpha_{i}-\beta_{j-1}}{\alpha_{i}-\alpha_{j}}=0.$ (69)
Proof.
Note that the LHS of Equation 69 can be written as
$\sum_{i=1}^{L}\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{i}-\beta_{j})\prod_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{L}\frac{1}{\alpha_{i}-\alpha_{j}}.$
Then we multiply $\prod_{j=2}^{L}(\alpha_{1}-\alpha_{j})$ on it and obtain
$\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{1}-\beta_{j})-\sum_{i=2}^{L}\prod_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{L}(\alpha_{1}-\alpha_{j})\cdot\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{i}-\beta_{j})\prod_{\begin{subarray}{c}j=2\\ j\neq i\end{subarray}}^{L}\frac{1}{\alpha_{i}-\alpha_{j}}.$ (70)
Then, we consider it as a polynomial in variable $\alpha_{1}$ with parameters $\alpha_{2},\ldots,\alpha_{L},\beta_{1},\ldots,\beta_{L-1}$. This polynomial is of degree $L-2$. Then, if we set $\alpha_{1}=\alpha_{k}$ for a $2\leq k\leq L$, we can see that Equation 70 becomes
$\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{k}-\beta_{j})-\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{L}(\alpha_{k}-\alpha_{j})\cdot\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{k}-\beta_{j})\prod_{\begin{subarray}{c}j=2\\ j\neq k\end{subarray}}^{L}\frac{1}{\alpha_{k}-\alpha_{j}}$
$=$ $\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{k}-\beta_{j})-\prod_{\begin{subarray}{c}j=1\\ j\neq m\end{subarray}}^{L-1}(\alpha_{k}-\beta_{j})$
[…]
Therefore, this polynomial has $L-1$ roots $\alpha_{2},\ldots,\alpha_{L}$, which means it must be the 00 polynomial. ∎
4.3 Auxiliary results
Proposition 4.7.
Let $C_{U}$ be that defined in Definition 3.4. We have
$\mathop{\bf E\/}_{U}[C_{U}]$ $=\bigoplus_{\lambda\vdash n+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|.$
Proof.
Note that $C_{U}=|U\rangle\!\rangle\!\langle\!\langle U|^{\otimes n+1}$ is an $(n+1)$-comb on $(\mathcal{H}_{1},\mathcal{H}_{2},\ldots,\mathcal{H}_{2n+2})$. We can easily check the following facts.
Fact 4.8.
$\mathop{\bf E\/}_{U}[C_{U}]$ commutes with the action of the group $\mathbb{U}_{d}\times\mathbb{U}_{d}$, i.e., for $(V,W)\in\mathbb{U}_{d}\times\mathbb{U}_{d}$:
$(V\otimes W)^{\otimes n+1}\mathop{\bf E\/}_{U}[C_{U}]=\mathop{\bf E\/}_{U}[C_{U}](V\otimes W)^{\otimes n+1}.$
Proof of 4.8.
$(V\otimes W)^{\otimes n+1}\mathop{\bf E\/}_{U}[C_{U}]$ $=\mathop{\bf E\/}_{U}\!\left[(V\otimes W|U\rangle\!\rangle\!\langle\!\langle U|)^{\otimes n+1}\right]$
$=\mathop{\bf E\/}_{U}\!\left[|VUW^{\textup{T}}\rangle\!\rangle\!\langle\!\langle U|^{\otimes n+1}\right]$
$=\mathop{\bf E\/}_{U^{\prime}}\!\left[|U^{\prime}\rangle\!\rangle\!\langle\!\langle V^{\dagger}U^{\prime}W^{*}|^{\otimes n+1}\right]$ (71)
$=\mathop{\bf E\/}_{U^{\prime}}\!\left[(|U^{\prime}\rangle\!\rangle\!\langle\!\langle U^{\prime}|V\otimes W)^{\otimes n+1}\right]$
$=\mathop{\bf E\/}_{U}[C_{U}](V\otimes W)^{\otimes n+1},$
where Equation 71 is because $U$ and $U^{\prime}$ are Haar random unitary. ∎
Fact 4.9.
$\mathop{\bf E\/}_{U}[C_{U}]$ is invariant under simultaneous permutation, i.e., for any $\pi\in\mathfrak{S}_{n+1}$, let $P^{\textup{odd}}(\pi)$ be the action of $\pi$ on $\mathcal{H}_{1}\otimes\mathcal{H}_{3}\otimes\cdots\mathcal{H}_{2n+1}\cong(\mathbb{C}^{d})^{\otimes n+1}$ and $P^{\textup{even}}(\pi)$ be that on $\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\mathcal{H}_{2n+2}\cong(\mathbb{C}^{d})^{\otimes n+1}$, then
$P^{\textup{even}}(\pi)\otimes P^{\textup{odd}}(\pi)\mathop{\bf E\/}_{U}[C_{U}]=\mathop{\bf E\/}_{U}[C_{U}]P^{\textup{even}}(\pi)\otimes P^{\textup{odd}}(\pi)=\mathop{\bf E\/}_{U}[C_{U}].$
Proof of 4.9.
$P^{\textup{even}}(\pi)\otimes P^{\textup{odd}}(\pi)\mathop{\bf E\/}_{U}[C_{U}]=\mathop{\bf E\/}_{U}\!\left[P^{\textup{even}}(\pi)\otimes P^{\textup{odd}}(\pi)|U\rangle\!\rangle\!\langle\!\langle U|^{\otimes n+1}\right]=\mathop{\bf E\/}_{U}\!\left[|U\rangle\!\rangle\!\langle\!\langle U|^{\otimes n+1}\right]=\mathop{\bf E\/}_{U}[C_{U}].$
The other direction is similar. ∎
Note that we have the following decomposition of the space
$(\mathcal{H}_{1}\otimes\mathcal{H}_{3}\otimes\cdots\otimes\mathcal{H}_{2n+1})\otimes(\mathcal{H}_{2}\otimes\mathcal{H}_{4}\otimes\cdots\otimes\mathcal{H}_{2n+2})\cong$ $\,\,(\mathbb{C}^{d})^{\otimes n+1}\otimes(\mathbb{C}^{d})^{\otimes n+1}$
$\stackrel{{\scriptstyle\mathbb{U}_{d}\times\mathbb{U}_{d}\times\mathfrak{S}_{n+1}\times\mathfrak{S}_{n+1}}}{{\cong}}$ $\bigoplus_{\lambda,\mu\vdash n+1}\mathcal{Q}_{\lambda}\otimes\mathcal{Q}_{\mu}\otimes\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu}.$
Therefore, by 4.8 and Schur’s lemma, we have
$\mathop{\bf E\/}_{U}[C_{U}]=\bigoplus_{\lambda,\mu\vdash n+1}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes M_{\lambda,\mu},$
for some $M_{\lambda,\mu}\in\mathcal{L}(\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu})$.
On the other hand, letting $P_{\mathop{\textup{avg}\/}}=\frac{1}{(n+1)!}\sum_{\pi\in\mathfrak{S}_{n+1}}P^{\textup{even}}(\pi)\otimes P^{\textup{odd}}(\pi)$, we have
$P_{\mathop{\textup{avg}\/}}$ $=\bigoplus_{\lambda,\mu\vdash n+1}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\frac{1}{(n+1)!}\sum_{\pi\in\mathfrak{S}_{n+1}}P_{\lambda}(\pi)\otimes P_{\mu}(\pi)$ (72)
$=\bigoplus_{\lambda,\mu\vdash n+1}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\mu}}\otimes\frac{1}{(n+1)!}\sum_{\pi\in\mathfrak{S}_{n+1}}P_{\lambda}(\pi)\otimes P^{*}_{\mu}(\pi)$ (73)
$=\bigoplus_{\lambda\vdash n+1}I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes\frac{1}{\dim(\mathcal{P}_{\lambda})}|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|,$ (74)
where $P_{\lambda}(\pi)$ is the action of $\pi$ on $\mathcal{P}_{\lambda}$, Equation 73 is because $P_{\mu}(\pi)$ is a real-valued matrix on the Young basis, Equation 74 is because the subspace in $\mathcal{P}_{\lambda}\otimes\mathcal{P}_{\mu}$ that is invariant under $P_{\lambda}(\pi)\otimes P_{\mu}^{*}(\pi)$ is $\textup{span}(|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle)$ when $\lambda=\mu$ and the trivial subspace $\{0\}$ when $\lambda\neq\mu$.
Then, 4.9 implies that $P_{\mathop{\textup{avg}\/}}\mathop{\bf E\/}_{U}[C_{U}]P_{\mathop{\textup{avg}\/}}=\mathop{\bf E\/}_{U}[C_{U}]$, which means
$\mathop{\bf E\/}_{U}[C_{U}]=\bigoplus_{\lambda\vdash n+1}z_{\lambda}\cdot I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|,$
for some $z_{\lambda}\in\mathbb{C}$.
Moreover, since $\operatorname{tr}_{2,4,\ldots,2n+2}(C_{U})=I^{\otimes n+1}$, we have
$I^{\otimes n+1}$ $=\bigoplus_{\lambda\vdash n+1}z_{\lambda}\cdot\operatorname{tr}_{\mathcal{Q}_{\lambda}}(I_{\mathcal{Q}_{\lambda}})\cdot I_{\mathcal{Q}_{\lambda}}\otimes\operatorname{tr}_{\mathcal{P}_{\lambda}}(|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|)$ (75)
$=\bigoplus_{\lambda\vdash n+1}z_{\lambda}\cdot\dim(\mathcal{Q}_{\lambda})\cdot I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{P}_{\lambda}},$ (76)
which means $z_{\lambda}=\frac{1}{\dim(\mathcal{Q}_{\lambda})}$. Therefore,
$\mathop{\bf E\/}_{U}[C_{U}]=\bigoplus_{\lambda\vdash n+1}\frac{1}{\dim(\mathcal{Q}_{\lambda})}\cdot I_{\mathcal{Q}_{\lambda}}\otimes I_{\mathcal{Q}_{\lambda}}\otimes|I_{\mathcal{P}_{\lambda}}\rangle\!\rangle\!\langle\!\langle I_{\mathcal{P}_{\lambda}}|.$
Lemma 4.10.
For any positive semidefinite matrix $M$ and vector $\lvert\psi\rangle$ such that $\lvert\psi\rangle\in\operatorname{supp}(M)$, we have
$M\sqsupseteq\lvert\psi\rangle\!\langle\psi\rvert\Longleftrightarrow 1\geq\langle\psi\rvert M^{-1}\lvert\psi\rangle,$
where $M^{-1}$ is the pseudo-inverse of $M$.
Proof.
$M\sqsupseteq\lvert\psi\rangle\!\langle\psi\rvert\Longleftrightarrow I_{\operatorname{supp}(M)}\sqsupseteq M^{-1/2}\lvert\psi\rangle\!\langle\psi\rvert M^{-1/2},$
where $M^{-1/2}$ is the pseudo-inverse of $M^{1/2}$. Then
$I_{\operatorname{supp}(M)}\sqsupseteq M^{-1/2}\lvert\psi\rangle\!\langle\psi\rvert M^{-1/2}\Longleftrightarrow 1\geq\operatorname{tr}(M^{-1/2}\lvert\psi\rangle\!\langle\psi\rvert M^{-1/2})=\langle\psi\rvert M^{-1}\lvert\psi\rangle.$
5 Hardness of unitary controlization
In Section 3, we have already seen the hardness of the unitary time-reversal task. In this section, we consider the closely related unitary controlization task: adding quantum control to an unknown unitary. More precisely, the task is formally stated in the following definition.
Definition 5.1 (Phase-oblivious unitary controlization).
Given queries to an unknown unitary $U$ of dimension $d$, the task is to implement $m$ queries to control-$U$ to within diamond error $\epsilon$; in particular, to implement $m$ unitaries $V_{1},\ldots,V_{m}$ such that
$\sum_{j=1}^{m}\left\lVert V_{j}-\left\lparen\lvert 0\rangle\!\langle 0\rvert\otimes I+e^{i\theta}\lvert 1\rangle\!\langle 1\rvert\otimes U\right\rparen\right\rVert_{\diamond}\leq\epsilon,$
for some $\theta\in[0,2\pi)$.
Note that in Definition 5.1, the task is oblivious to the relative phase of the control-$U$: the angle $\theta$ in the relative phase $e^{i\theta}$ can be arbitrary and not specified by the task. We only require that all $V_{1},\ldots,V_{m}$ approximately implement the same control-$U$ up to the same relative phase. This phase-oblivious requirement is weaker than those in the previous literature of quantum unitary controlization [Soe13, AFCB14, TMVG18, GST24, TMM+25, DNSM19] (see also [ZRK+11, FDDB14, BDP16]), where the relative phase is specified by the task. Actually, in the latter case, it is impossible to solve the unitary controlization task using a finite number of queries [GST24]. In context without ambiguity, we will simply call Definition 5.1 the unitary controlization task.
The main theorem of this section is a query lower bound for the unitary controlization task in Definition 5.1, as follows.
Theorem 5.2 (Hardness of unitary controlization).
Any protocol that solves the unitary controlization task in Definition 5.1, even for constant $m$ and $\epsilon$, must us at least $\Omega\left\lparen d^{2}\right\rparen$ ) queries to the unknown $d$-dimensional unitary $U$.
5.1 Time-reversal from controlled unitary
To prove Theorem 5.2, we will need the following Lemma 5.3 on approximately reversing an unknown unitary $U$ using queries to control-$U$, based on a folklore use of the coherent phase estimation [CEMM98, NC10]. Equipped with Lemma 5.3 and the previous Corollary 3.2, we are able to prove Theorem 5.2, as shown later in this section.
Lemma 5.3 (Time-reversal from controlled unitary).
Given queries to control-$U$ for some unknown unitary $U$, there is a protocol to implement the time-reverse $U^{-1}$ to within diamond norm error $\epsilon$ using $O(1/\epsilon^{3})$ queries.
In Lemma 5.3, we allow the control-$U$ to be any unitary of the form $\lvert 0\rangle\!\langle 0\rvert\otimes I+e^{i\theta}\lvert 1\rangle\!\langle 1\rvert\otimes U$, where $\theta\in[0,2\pi)$ can be unknown. Since the unitary $U$ and its time-reverse $U^{-1}$ are oblivious to $\theta$, we can prove Lemma 5.3 just for the case $\theta=0$, without loss of generality. To prove Lemma 5.3, let us first recall the standard coherent phase estimation [CEMM98]. While there are algorithms for this task with better efficiency [NWZ09, PTSS24], here we only consider the following lemma adapted from the textbook [NC10], for simplicity of later analysis. The choice does not affect our final Theorem 5.2.
Lemma 5.4 (Coherent phase estimation [CEMM98, NC10]).
Suppose that $U=\sum_{j=0}^{d-1}e^{2\pi i\varphi_{j}}\lvert v_{j}\rangle\!\langle v_{j}\rvert$ is any unknown unitary. There is a protocol to implement the coherent phase estimation of $U$ on any unknown input state with error $\epsilon$, using $O\left\lparen\frac{1}{\epsilon^{3}}\right\rparen$ ) queries to control-$U$. Specifically, let $a$ be the number of ancilla qubits for storing the estimated phases. Then the protocol implements a unitary $V$ satisfying
$V\lvert v_{j}\rangle\lvert 0\rangle_{a}=\lvert v_{j}\rangle\lvert\Phi_{j}\rangle_{a}$ (77)
for some $a=O\left\lparen\log(1/\epsilon)\right\rparen$ ) and $j=0$ to $d-1$, where $\left\lVert\Pi_{j}\lvert\Phi_{j}\rangle\right\rVert^{2}\geq 1-\epsilon^{2}$ and $\Pi_{j}=\sum_{\left\lvert\varphi_{j}-x_{j}/2^{a}\right\rvert\leq\epsilon}\lvert x_{j}\rangle\!\langle x_{j}\rvert$.
Using the coherent phase estimation in Lemma 5.4, we can (approximately) manipulate the eigenvalues of the unknown $U$ by performing operations conditioned on the estimated phases. Meanwhile, it is important to uncompute the estimated phases afterwards. The following corollary describe how such uncomputation can also be performed in a similar way.
Corollary 5.5 (Coherent phase uncomputation).
There is also a protocol to uncompute the estimated phases in Equation 77, using $O\left\lparen\frac{1}{\epsilon^{3}}\right\rparen$ ) queries to control-$U$; that is, to implement a unitary $W$ such that
$W\lvert v_{j}\rangle\lvert\Phi_{j}\rangle_{a}=\lvert v_{j}\rangle\lvert 0\rangle_{a}$
for $j=0$ to $d-1$.
Proof.
To prove Corollary 5.5, let us recall how the standard textbook algorithm [NC10] for Lemma 5.4 is constructed. The unitary $V$ in Lemma 5.4 is implemented in the following way.
  1. 1.
    Apply $H^{\otimes a}$ gates to the ancilla system.
  2. 2.
    For $k=0$ to $a-1$, apply control-$U^{2^{k}}$ gate, with the $(k+1)^{\textup{th}}$ ancilla qubit as control and the principal system as target.
  3. 3.
    Apply an inverse quantum Fourier transform to the ancilla system.
For the above implementation, we can explicitly write out
$\lvert\Phi_{j}\rangle=\frac{1}{2^{a}}\sum_{x,y=0}^{2^{a}-1}e^{-2\pi ixy/2^{a}}e^{2\pi i\varphi_{j}x}\lvert y\rangle.$ (78)
Now to uncompute the estimated phases, we can implement $W$ as follows.
  1. 1.
    Apply an arithmetic circuit to the ancilla system performing the unitary mapping: $\lvert x\rangle_{a}\mapsto\lvert 2^{a}-x\rangle_{a}$.
  2. 2.
    Apply a quantum Fourier transform to the ancilla system.
  3. 3.
    For $k=0$ to $a-1$, apply control-$U^{2^{k}}$ gate, with the $(k+1)^{\textup{th}}$ ancilla qubit as control and the principal system as target.
  4. 4.
    Apply $H^{\otimes a}$ gates to the ancilla system.
To see that $W$ correctly uncomputes the estimated phases, it suffices to verify that
$W^{\dagger}\lvert v_{j}\rangle\lvert 0\rangle_{a}=\lvert v_{j}\rangle\lvert\Phi_{j}\rangle_{a}.$ (79)
Let $Q$ be the associated unitary for the last three steps of $W$. It is easy to see that $Q^{\dagger}$ is exactly the coherent phase estimation of $U^{\dagger}$. Suppose that $Q^{\dagger}\lvert v_{j}\rangle\lvert 0\rangle_{a}=\lvert v_{j}\rangle\lvert\Psi_{j}\rangle_{a}$. Similar to Equation 78, we can explicitly write out $\lvert\Psi_{j}\rangle$ as
$\lvert\Psi_{j}\rangle=\frac{1}{2^{a}}\sum_{x,y=0}^{2^{a}-1}e^{-2\pi ixy/2^{a}}e^{-2\pi i\varphi_{j}x}\lvert y\rangle=\frac{1}{2^{a}}\sum_{x,y=0}^{2^{a}-1}e^{-2\pi ixy/2^{a}}e^{2\pi i\varphi_{j}x}\lvert-y\rangle.$ (80)
Applying the unitary mapping $\lvert x\rangle\mapsto\lvert 2^{a}-x\rangle$ to the RHS of Equation 80 yields $\lvert v_{j}\rangle\lvert\Phi_{j}\rangle$, which verifies Equation 79. ∎
Using coherent phase estimation (Lemma 5.4) and its associated phase uncomputation (Corollary 5.5), we can prove Lemma 5.3 by designing a protocol to reverse the unknown unitary $U$.
Proof of Lemma 5.3.
Suppose $U=\sum_{j=0}^{d-1}e^{2\pi i\varphi_{j}}\lvert v_{j}\rangle\!\langle v_{j}\rvert$. To implement the time-reverse $U^{-1}$ to within diamond norm error $\epsilon$, it suffices to implement a unitary $U^{\prime}$ such that for any $j=0$ to $d-1$:
$\left\lVert U^{\prime}\lvert v_{j}\rangle-e^{-2\pi i\varphi_{j}}\lvert v_{j}\rangle\right\rVert\leq\epsilon/2.$ (81)
This is because of the following inequalities between diamond distance, trace distance, and Euclidean distance:
$\left\lVert U^{\prime}-U^{-1}\right\rVert_{\diamond}\leq 2\left\lVert U^{\prime}-U^{-1}\right\rVert_{\operatorname{tr}}=2\max_{\lvert v\rangle}d_{\operatorname{tr}}\left\lparen U^{\prime}\lvert v\rangle,U^{-1}\lvert v\rangle\right\rparen\leq 2\max_{j}\left\lVert U^{\prime}\lvert v_{j}\rangle-e^{-2\pi i\varphi_{j}}\lvert v_{j}\rangle\right\rVert.$
Suppose $\epsilon^{\prime}$ is a parameter to be determined. Let us construct $U^{\prime}=W(I\otimes E)V$, where $V$ and $W$ are unitaries for phase estimation and uncomputation with error $\epsilon^{\prime}$ in Lemmas 5.4 and 5.5, respectively, and $E$ performs the unitary mapping $\lvert x\rangle_{a}\mapsto e^{-2\pi ix/2^{a}}\lvert x\rangle_{a}$. It remains to prove that $U^{\prime}$ satisfies Equation 81. According to Lemma 5.4, we first have
$\left\lVert V\lvert v_{j}\rangle\lvert 0\rangle_{a}-\lvert v_{j}\rangle\lvert\Phi_{j}^{\prime}\rangle_{a}\right\rVert\leq\epsilon^{\prime},$
where $\lvert\Phi_{j}^{\prime}\rangle=\Pi_{j}\lvert\Phi_{j}\rangle$ is unnormalized, and $\Pi_{j}$ is defined in Lemma 5.4. Consequently,
$\left\lVert(I\otimes E)V\lvert v_{j}\rangle\lvert 0\rangle_{a}-e^{-2\pi i\varphi_{j}}\lvert v_{j}\rangle\lvert\Phi_{j}^{\prime}\rangle_{a}\right\rVert\leq(2\pi+1)\epsilon^{\prime},$ (82)
since $\left\lvert e^{-2\pi i\varphi_{j}}-e^{-2\pi ix/2^{a}}\right\rvert\leq 2\pi\epsilon^{\prime}$ for any $x$ with $\left\lvert\varphi_{j}-x/2^{a}\right\rvert\leq\epsilon^{\prime}$. Finally, we can obtain
$\left\lVert W(I\otimes E)V\lvert v_{j}\rangle\lvert 0\rangle_{a}-e^{-2\pi i\varphi_{j}}\lvert v_{j}\rangle\lvert 0\rangle_{a}\right\rVert\leq(2\pi+2)\epsilon^{\prime},$
because $\left\lVert\lvert\Phi_{j}\rangle-\lvert\Phi^{\prime}_{j}\rangle\right\rVert\leq\epsilon^{\prime}$. By taking $\epsilon^{\prime}=O(\epsilon)$, the conclusion of the theorem immediately follows. ∎
5.2 Proof of Theorem 5.2
Finally, we are ready to prove the main Theorem 5.2 of this section.
Proof of Theorem 5.2.
Assume for contradiction that there exists a protocol that solves the unitary controlization task in Definition 5.1 using $n=o(d^{2})$ queries to $U$. Specifically, for any $m=O(1)$ and $\epsilon=O(1)$, we can implement $m$ unitaries $V_{1},\ldots,V_{m}$, using $o(d^{2})$ queries to $U$, such that
$\sum_{j=1}^{m}\left\lVert V_{j}-\left\lparen\lvert 0\rangle\!\langle 0\rvert\otimes I+e^{i\theta}\lvert 1\rangle\!\langle 1\rvert\otimes U\right\rparen\right\rVert_{\diamond}\leq\epsilon,$
for some $\theta\in[0,2\pi)$. By Lemma 5.3, there is a protocol to implement $U^{-1}$ to within diamond norm error $\epsilon$ using $l=O(1)$ queries to control-$U$. Let $m=l$, and we replace the $j^{\textup{th}}$ query to control-$U$ in this protocol by the unitary $V_{j}$ for $j=1$ to $m$. From the triangle inequality, we obtain a new protocol to implement $U^{-1}$ to within diamond norm error $2\epsilon$, using $nm=o(d^{2})$ queries to $U$. However, this is a contradiction with Corollary 3.2, and the conclusion immediately follows. ∎